### v1

# Cohort Discovery: A Survey on LLM-Assisted Clinical Trial Recruitment

Abstract
Recent advances in LLMs have greatly improved general-domain NLP tasks. Yet, their adoption in critical domains, such as clinical trial recruitment, remains limited. As trials are designed in natural language and patient data is represented as both structured and unstructured text, the task of matching trials and patients benefits from knowledge aggregation and reasoning abilities of LLMs. Classical approaches are trial-specific and LLMs with their ability to consolidate distributed knowledge hold the potential to build a more general solution. Yet recent applications of LLM-assisted methods rely on proprietary models and weak evaluation benchmarks. In this survey, we are the first to analyze the task of trial-patient matching and contextualize emerging LLM-based approaches in clinical trial recruitment. We critically examine existing benchmarks, approaches and evaluation frameworks, the challenges to adopting LLM technologies in clinical research and exciting future directions.
Cohort Discovery: A Survey on LLM-Assisted Clinical Trial Recruitment
Shrestha Ghosh1   Moritz Schneider2   Carina Reinicke2   Carsten Eickhoff1 1University of Tübingen, Germany 2Boehringer Ingelheim, Germany 1{first.last}@uni-tuebingen.de   2{first.last}@boehringer-ingelheim.com
1 Introduction
Clinical trials evaluate the effects of an intervention on human health. Selecting the precise and required size of patient population is crucial for trial completion. According to various estimates, more than 50% of aborted clinical trials fail due to low accrual rates, and 80% of all clinical trials do not manage to recruit the required patient cohorts within the allotted time Clinical Trials Arena (2012); Williams et al. (2015); Pharmaceutical Technology (2019). While the trend has steadily declined over the past decade, with the intensive use of technology-aided solutions, efficient patient recruitment remains the most crucial bottleneck in clinical trial research Clinical Trials Arena (2022), thereby hampering the rate at which potentially life-saving therapies or new vaccines can be introduced.
Figure 1: Steps in a typical patient recruitment process vs. our proposed flow with expert interaction.
Clinical trials have a dual nature, consisting of both trial-agnostic and trial-specific requirements, making it challenging to design generalized approaches. This has traditionally resulted in highly specialized matching processes that work for a particular trial or patient population. The resulting patient recruitment process is linear, as illustrated (in orange) in Figure 1, leaving little room for expert intervention and feedback. It starts with extracting information from a trial description and available patient records, then matching explicit trial eligibility criteria and patient data to determine patient eligibility. Finally, experts review the selected patients before contacting them.
An expert reviews several hundred patients Ni et al. (2015) and can end up spending hours on one patient, incurring significant costs Penberthy et al. (2012). In their study on five trials, Penberthy et al. (2010) showed that automated matching, via table queries and string-matching algorithms, saved between 165 hours to 1,329 hours of reviewing time when compared to manual evaluation. As more patient data becomes digitally available, research into applying machine intelligence and developing explainable systems for clinical researchers to correctly interpret model predictions has gained traction Murdoch and Detsky (2013); Payrovnaziri et al. (2020); von Itzstein et al. (2021).
Generative LLMs can serve as knowledge aggregators and streamline clinical trial design and trial-patient matching (in purple in Figure 1). However, the clinical trial recruitment process has seen relatively low adoption of transformer language models He et al. (2020); Harrer (2023); Lu et al. (2024). Instruction fine-tuned LLMs for generating explainable trial-patient matches has received attention only since the past one year Jin et al. (2024); Nievas et al. (2024); Rybinski et al. (2024); Wornow et al. (2025). Given the sensitive nature of clinical trials and the rapidly evolving landscape of LLM technology, there is no prior work surveying this area. We fill this gap by critically examining the trial-patient matching approaches, the method used, the benchmarks available and the evaluation strategies. The existing systematic reviews of NLP methods and their performance Kim and Quintana (2022); Idnay et al. (2022); Kantor and Morzy (2024) are desiged to follow strict inclusion/exclusion criteria and search strategies Khan et al. (2003); Moher et al. (2009). This practice, essential in medical research for measuring the association between intervention and outcomes, misses the larger picture of contextualizing LLMs in clinical trial recruitment settings. In contrast, this work follows the perspective of surveys in computer science (see Appendix B) to provide a more comprehensive view of the role of LLMs in clinical trial recruitment.
Our Contribution
Cohort discovery has been a long-standing problem and LLM-assisted methods to improve trial-patient matching has only recently gained momentum. In this work, we formalize the problem of trial-patient matching, and differentiate between two major approaches (trial-centric and patient-centric), the associated tasks (classification vs. ranking) and their effect on the choice of benchmarks (longitudinal vs. short patient descriptions, single vs. multi-trial). We identify the important bottlenecks in data availability and task formalization. We introduce a taxonomy of errors in LLM-generations for trial-patient matching. Finally, we provide actionable steps for the community on four promising directions.
2 Background
Clinical Trial Recruitment
Also known as patient recruitment/enrollment/(pre-)screening, it is the process of finding relevant patients (or cohort) for a clinical trial. The steps in patient recruitment, as shown in Figure 1, include data sourcing, information extraction, matching, and expert review.
Eligibility Criteria
They are explicit conditions for (inclusion) or against (exclusion) enrollment, usually expressed in natural language and designed by trial investigators to select cohorts. These conditions have varying difficulty levels, ranging from simple tabular operations (age or gender requirements) to semantic matching (medical condition or medications) and inference (reasoning with negation, patient history).
Patient Records
They refer to any form of data, in a combination of structured and unstructured formats, used to describe a patient. When digitized, it is often called an Electronic Health Record (EHR). This data includes the patient’s demographic information (age, gender, ethnicity), clinical notes, medications, billing history, and test results, that are collected over a period of years.
Biomedical NLP and LLMs
Transformer language models have greatly improved information retrieval and extraction tasks. There exist specialized resources, such as, PubmedQA Jin et al. (2019), MIMIC-III & IV Johnson et al. (2016, 2020) and BioRED Luo et al. (2022), to train encoder language models Radford (2018); Devlin et al. (2019); Raffel et al. (2020) on medical domain (e.g., BioBERT Lee et al. (2020), PubmedBERT Gu et al. (2021), Criteria2Query2.0 Fang et al. (2022), MedCPT Jin et al. (2023)). Yet, Dorfner et al. (2024) show that fine-tuning generative LLMs on biomedical domain has limited relative performance gain. In fact, Alber et al. (2025) show that medical misinformation encountered during the fine-tuning affect downstream tasks. The Journal of American Medical Informatics Association (JAMIA), which published 41 articles on biomedical health and LLMs in a focus issue, observed a shift from domain-specific small models towards larger general-purpose LLMs Lu et al. (2024). Even among LLMs, the OpenAI Achiam et al. (2023) family of proprietary models (GPT3.5 and GPT4 being the most common) is used far more often than open-sourced LLaMA Touvron et al. (2023) or Mistral Jiang et al. (2023) models. Finally, Bedi et al. (2025) report that of more than 519 biomedical studies using LLMs, less than 5% used real patient data, tasks like summarization are understudied and fairness, bias, uncertainty and deployment considerations are rarely measured.
State of Adoption of NLP Advancements
Historically, patient recruitment has been a manual process with minimal automation via pre-defined database queries Thadani et al. (2009), and, more recently, automatic translations of free text into database queries Yuan et al. (2019); Fang et al. (2022); Park et al. (2024). A recent review reported low adoption rates of AI, particularly NLP, in clinical trials Kantor and Morzy (2024). Another review on understanding the role of NLP systems in patient recruitment found just 11 studies that evaluated NLP systems in clinical settings Idnay et al. (2022). Heterogeneous outcomes, diverse results and a dependence on small retrospective data drive the gap in NLP research and their adoption in real-world settings Idnay et al. (2024). A study by Idnay et al. (2023), investigating how clinical researchers screened patients, highlights the challenges of universal and domain-specific nature of the eligibility criteria and makes recommendations to build interactive, flexible and transparent recruitment strategies. Interestingly, when Corbaux et al. (2024) categorize trial matching tools for oncological trials as manual, semi-automatic and automatic, automatic methods still fall in the research and development phase, yet to be commercially available.
Review Methodology
We searched for relavant papers since 2019, using the keywords “clinical trial”, “cohort discovery”, “patient recruitment”, “trial recruitment”, “trial matching” in conjunction with “llm”, “language model”, “gpt”. We included method and dataset papers using LLMs and excluded papers that used LLMs in biomedical NLP (broader scope) or other health care applications (not clinical trials). Starting with the citations in the LLM papers we tracked the pre-LLM literature by recursively tracing their citations in newer publications. Since the topic “clinical trial recruitment” is vast, this approach efficiently captured the evolution of the predictive methods in this field.
3 Trial-Patient Matching
Trial-patient matching assigns a trial-patient pair one of the three labels of eligible, ineligible and irrelevant based on the patient record and eligibility criteria. Importantly, a trial-centric approach, taken by a trial investigator, matches patients from a hospital database to a specific trial. Alternatively, a patient-centric approach, taken by a patient or their healthcare provider, matches relevant trials from a trial registry to a patient.
Given sets of inclusion and exclusion criteria ($C_{inc},C_{exc}$) from a trial and a set of patient data, $P$, we formalize the trial-patient matching problem, $M$, to predict one of the labels Ineligible (Inel.), Irrelevant (Irr.) or Eligible (Eli.) by aggregating criterion-level binary matches $M^{\prime}(c,P)$.
$\!M(C_{inc},C_{exc},P)\!=\!\left\{\!\begin{array}[]{ll}\!\textit{Inel.},\!&\!\exists c\in C_{exc},M^{\prime}(c,P)\\ \!\textit{Irr.},\!&\!\exists c\in C_{inc},\neg M^{\prime}(c,P)\\ \!\textit{Eli.},\!&\!\neg(\textit{Inel.}\lor\textit{Irr.})\end{array}\!\right.$
This induces a priority, such that, a patient satisfying any exclusion criteria becomes ineligible, regardless of inclusion criteria matches. In Section 6, we elaborate on the lack of a formal task definition and its effects on trial-centric matching.
Task Benchmark Size Eval. Metrics Best Score Cohort
selection
2018 N2C2
Stubbs et al. (2019)
#patients = 288
#records per patient = 2-5
#tokens/patient = 2711
#trials = 1
#criteria = 13
Micro-averaged:
P, R, F1
0.91 micro F1
Trial
ranking
Koopman and Zuccon (2016) #patient summaries (22 words avg) = 60
#patient description (78 words avg) = 60
#keywords (4.4 words avg) = 489
#trials = 204,855
#relevant matches = 685
#relevance judgements = 4000
MRR,
P@5,
adaptive precision
0.3 MRR
$<$ 0.2 P@5
TREC CT Track 2021
Soboroff (2021)
#patient descriptions = 75
#trials = 375,580
#relevant matches = 5,570
#relevance judgements = 35,832
NDCG@10,
P@10,
MRR,
R-Precision
0.84 NDCG@10
0.748 P@10
1.0 MRR
TREC CT Track 2022
Roberts et al. (2022)
#patient descriptions = 50
trials = 375,580
#relevant matches = 3,949
#relevance judgements = 35,394
NDCG@10,
P@10,
MRR,
R-Precision
0.61 NDCG@10
0.50 P@10
0.32 R-Precision
0.72 MRR
TREC CT Track 2023
Rybinski et al. (2024)∗∗
#patient tables = 40
#disease templates = 8
#trials = 451,538
#relevant matches = 11,963
#relevance judgements = 34,931
NDCG@10,
P@10,
MRR
0.81 NDCG@10
0.73 P@10
0.78 MRR
Best scores are aggregated from the Appendix of the proceedings of TREC 2021. ∗∗Borrowed from Rybinski et al. (2024) as TREC CT 2023 does not have a published track overview.
Table 1: Overview of public evaluation benchmarks for trial-patient matches covered in this survey.
4 Benchmarks and Corpora
We analyze five public benchmarks in context of the corpora and their evaluation metrics.
  • 2018 N2C2 Cohort Selection for Clinical Trials Stubbs et al. (2019). [^1] Currently unavailable as of 2024 Nov 6.
  • Trial ranking Koopman and Zuccon (2016).
  • Text REtrieval Conference Clinical Trial (TREC CT) [^2] Available at https://www.trec-cds.org/. tracks 2021, 2022, 2023 for ranking trials.
Table 1 provides an overview of the benchmarks.
4.1 Trial Data
ClinicalTrials.gov is one of the largest online database of clinical studies, including clinical trials and other observational studies. As of November 2024, it lists more than 390,000 clinical trials, thousands of which are active. These studies are submitted by investigators from over 200 countries. The euclinicaltrials.eu is another online registry for trials conducted in the European Union. It comprises more than 50,000 trials, with more than 7000 being active. Every trial comprises a study description, eligibility criteria and study plan among other details. The ClinicalTrials.gov database forms the primary source of trials in many academic studies.
As we will see in Section 5, earlier trial-specific approaches manually separated discrete data from free-text search. However, this separation is difficult to scale, and newer approaches focus on semantic matching. The N2C2 cohort selection benchmark is based on a single trial. Hence, systems specifically model the trial criteria. The TREC CT benchmarks source the ClinicalTrials.gov database for the task of trial ranking. However, additional annotation of the eligibility criteria denoting their type or complexity is unavailable.
4.2 Patient Data
Since patient data contains sensitive information, additional measures are required to remove personal information in order to release them as public repositories. The MIMIC database Johnson et al. (2016, 2020) is the largest public database of patient information. The latest version, MIMIC-IV, contains rich data on hospital admissions, medications, demographic data and notes of 299,712 de-identified patients. This dataset is not used in any of the five benchmarks. One reason might be a significant challenge in obtaining confirmed trial-patient match labels, which is easily available retrospectively from existing hospital records. The N2C2 cohort selection benchmark Stubbs et al. (2019), contains 288 de-identified longitudinal records of patients. The TREC CT benchmarks use synthetic patient profiles, comprising short textual patient descriptions and keywords defining cohorts.
4.3 Evaluation Metrics
Standard IR metrics (precision, recall, F1) and their derivatives (precision at rank (P@k), normalized discounted cumulative gains (NDCG@k), R-precision, mean reciprocal rank (MRR) and adaptive precision Moffat et al. (2015)) are the most popular evaluation metrics. Accuracy, specificity, and workload are also prevelant in the medical community. Notably, researchers in the medical domain have stressed the lack of definition and consistency of ground truth data to evaluate NLP tools in clinical trial matching Kim and Quintana (2022). The trial-patient ground-truth labels are obtained manually either from medical experts Koopman and Zuccon (2016); Soboroff (2021); Roberts et al. (2022) or retrospectively from past trials Stubbs et al. (2019). The TREC CT annotations were created from the top-k results pooled among all participating teams, resulting in tens of thousands of relevance judgments. Out of a total of 35,394 pooled trial-patient matches in the 2022 edition, 11% were judged as eligible, 9% as ineligible and 80% as not relevant with an average of 700 trials judged per patient Roberts et al. (2022). In the 2023 edition, the relevance judgments were more balanced between the three labels Rybinski et al. (2024). Measurement of human preference of LLM generations, bias in model predictions and the informativeness of LLM explanations ranges from limited to non-existent in the current works.
5 Trial-Patient Matching Methods
We broadly categorize systems into classical and LLM-based approaches and further group them by their direction of approach, trial/patient-centric. The LLM-based approaches focus on tasks, such as, criterion/trial-level predictions and relevance ranking, which in principle is direction-agnostic. Appendix A lists the works cited in this paper.
5.1 Classical Approaches
Trial-centric Matches
Majority of classical models use this approach build trial or cohort specific models. They follow a two-step method, starting with manually-designed rule-based discrete data filters, followed by sub-word matching Penberthy et al. (2010), vector matching Ni et al. (2015) or machine-learned classifiers Tun et al. (2023) for matching eligibility criteria or cohort keywords. Ni et al. (2015) showed a retrospective workload reduction to 50 patients down from 163 per trial. IBM Watson for clinical trial matching (IBM-CTM) is a proprietary commercial tool for trial-patient matching, which uses neural NLP techniques to match eligibility criteria to patients Helgeson et al. (2018). While one study on 3 trials measured a reduction in workload from 110 minutes to 86 minutes for screening 90 patients Beck et al. (2017), another study on 10 clinical trials and 102 patients evaluated the agreement in matches between IBM-CTM and expert annotation to be above 80% Alexander et al. (2020). Another line of work that focuses on defining a cohort by translating natural language criteria into formal query language for instantly querying patient databases Yuan et al. (2019).
Patient-centric Matches
Ni et al. (2015) create bag-of-words vectors to retrieve top trial matches for a patient and show a retrospective workload reduction (4 vs. 42 trials per patient) and 33% precision on a subset of 127 patients. The workload increases to 20 trials per patient and the precision drops to less than 4% when generalized to 215 patients. Zhang and Demner-Fushman (2017) train an SVM classifier on the top unigram and bigram features to predict if a trial is eligible for a cohort (e.g., HIV patients). Additional features derived from NER and lemmatization yielded small gains in performance.
5.2 LLM-based Approaches
These approaches do not rely on trial-specific heuristics. Yet, as systems are evaluated on a few disease-specific trials, the largest being 146 trials, covering 10 cancer types, evaluated by Hamer et al. (2023), the generalizability of LLMs across diseases is unclear. Moreover, when compared, LLMs are less effective in making trial-level predictions (low F1 scores) than criterion-level ones.
Trial-Level Prediction
Wong et al. (2023) used GPT4 to structure eligibility criteria into logical expressions. Despite lagging behind expert systems, which use heavy manual intervention, they demonstrate the readiness of general-purpose LLMs for cohort-specific NER (oncology in this case) and logical structuring. Zero-shot and few-shot LLM prompts outperformed the state-of-the-art biomedical NER, SciSpaCy Neumann et al. (2019).
Yuan et al. (2023) prompt LLMs to augment eligibility criteria and pass the BERT-based embeddings of these criteria and patient data through a fully connected classification layer to predict patient-criterion eligibility. The trial-level prediction is then composed by asserting that all inclusion criteria and none of the exclusion criteria match. On six trials for stoke patients, the model achieved 0.91 F1 at criterion-level vs. 0.81 F1 at trial-level.
Criterion-Level Prediction
Hamer et al. (2023) use OpenAI’s text-davince-003 model to first filter criteria that applies to a patient, then justify whether applicable criteria are met and to finally predict criterion-level labels based on previous explanations. This effectively reduces the total number of eligibility criteria (down to 300 from originally 4,100) a physician has to assess before recommending trials to a patient.
Wornow et al. (2025) prompt LLMs with trial-patient pairs to generate structured JSON strings of criterion-level predictions, explanations with evidence and a self-declared confidence. The content of the prompt ranges from the entire patient record and a trial to a single criterion and relevant record snippets. GPT4 achieves a micro-F1 of 0.93 vs. 0.91 points by the prior state-of-the-art on the 2018 N2C2 cohort dataset. The RECTIFIER Unlu et al. (2024) employs a RAG-based approach to find top-k relevant patient data stored as vector embeddings. The criteria and relevant records are prompted to the LLM to generate either Yes or No. Evaluated on just a single trial and a pool of 3000 private patient records, with clinical notes spanning three years, RECTIFIER achieved an accuracy higher than trained non-licensed staff overall (92.7% vs. 89.1%), especially for the Symptomatic Heart Failure criterion (97.9% vs. 91.7%).
Trial Ranking
One line of work investigates the role of neural re-rankers for ranking relevant trials. Pradeep et al. (2022) generate multiple independent queries from a patient description to retrieve relevant trials. The trial and patient description is then passed to a fine-tuned T5 model to generate a relevance label. Zhuang et al. (2023) propose a multi-stage pipeline that involves a hybrid sparse-dense retriever, initialized with PubmedBERT-base and fine-tuned to retrieve the top 1000 trials. A cross-encoder, initialized with PubmedBERT-large, is trained to re-rank these results. Finally, the top-20 trials are re-ranked using GPT4. Rybinski et al. (2024) leverage LLMs for query reformulation and trial re-ranking to outperform traditional BM25 and BERT-based retrieval models. They show that BM25 is an effective initial retriever, surpassed by BERT-based neural re-rankers.
Jin et al. (2024) use GPT4 to first extract query keywords to retrieve top-k trials and then prompt GPT4 with patient data and eligibility criteria to generate criterion-level labels and explanations. They use a combination of manually-defined and GPT4 prompted scores to determine a trial-level score for final raning. Their TrialGPT performs better on the TREC datasets compared to smaller domain-specific transformer models and show a screening time reduction by 42%. Trial-LLama by Nievas et al. (2024) extend this study by extensive fine-tuning of open-source LLaMA models.
6 Discussion
Annotated Corpora
There have been few attempts to analyze and structure the clinical trials repository. As a result, finding similar trials and analyzing systematic failure cases is notoriously difficult. Chia, a manually annotated corpus of 1000 clinical trials created by Kury et al. (2020) has criterion-level annotations of entities, relations and Boolean logic graphs. Such annotations on trial-patient matching benchmarks can help us measure model fairness. Scaling patient data collections is also challenging as evident from Table 1, where patient records are in the hundreds. Individual records, curated manually, range from keywords via short patient summaries to documents over a time period. The format ranges from free text via semi-structured in-line tables to full-fledged databases. The rich and large MIMIC dataset, remains largely untapped for the task of trial-patient matching. LLMs provide a promising avenue for solving the automatic annotation at scale.
Grouth-truth Labels
Access to trial-patient eligibility labels, i.e., {eligible, ineligible, not-relevant}, is integral for training and evaluation. As they are obtained from human experts, they are difficult to scale and take time. Hence, many systems are evaluated retrospectively on patient data from private repositories such as clinics that maintain participation records in past trials Wong et al. (2023); Yuan et al. (2023); Hamer et al. (2023); Unlu et al. (2024). The largest public annotations from the TREC CT tracks have an average of 700 trial annotations per patient description. Bedi et al. (2025) found that of the 519 studies between 2022 and 2024 that evaluated LLM applications in healthcare, only 5% used real patient data. Jointly, the major public data sources, e.g., ClinicalTrials.gov and MIMIC datasets, present an exciting opportunity to develop generalizable matching methods.
Formalization
We formalize the task of trial-patient matching in Equation 3 through simple set operations. Surprisingly, except Yuan et al. (2023), none of the related work explicitly formalize the task. This coincides with a lack of aggregation strategies at the criterion-level Hamer et al. (2023); Wornow et al. (2025); Unlu et al. (2024). Formalizing the task, guides Yuan et al. (2023) to design a loss function that accounts for the contrastive requirements of inclusion and exclusion criteria. The criterion-level labels by Yuan et al. (2023), while not being directly interpretable, are aggregated as in Equation 3 for the trial-level labels. Wong et al. (2023) could potentially explain prediction via logical component mismatches, though the current F1 measures are below 50 points. Clinical experts, who analyze matches for the final cohort selection, strictly follow Equation 3. Criterion-level predictions and explanations have shown to reduce screening time by 42% Jin et al. (2024).
Core Limitations of LLMs
According to Harrer (2023) the core limitations affecting LLM adoption are unfiltered pre-training, which does not differentiate between facts, opinions, or misinformation; lack of self-assessment, where a model generates invalid but syntactically and semantically coherent sentences; non-determinism: where surface-form prompt variations lead to drastic changes in the output and repeatability is not guaranteed under consistent input conditions; knowledge recall: where updating outdated data or injecting new information requires expensive retraining since the mechanisms of memory in LLMs are not well understood. In Section 7, we discuss how knowledge recall and lack of self-assessment surface through reasoning errors. These limitations pose a direct challenge to the transparency and accountability principles of AI for health laid down by the World Health Organization Guidance (2021).
Societal Impact
Recent research measured distinct bias in disease diagnosis across gender, age and disease in popular LLM models (GPT4, ChatGPT and Qwen) Zhao et al. (2024b) and found that GPT4 tends to stereotype demographic presentations when generating diagnoses Zack et al. (2024). There is a lack of such analyses and mitigation methods in trial-patient matching and a lack of diverse datasets. The popularity of propreitary models for biomedical tasks Lu et al. (2024) including trial-patient matching Wornow et al. (2025); Jin et al. (2024) further ethical concerns on handling sensitive patient data. Alber et al. (2025) show that LLMs are prone to making medical misjudgments by replacing just 0.001% of the training data with medical misinformation. More research is required to determine the feasibility of locally deploying open-source models, the associated costs of maintaining them and keeping them up-to-date.
Figure 2: Taxonomy of errors in LLM-generations.
7 Taxonomy of LLM Errors
Since there is no one accepted taxonomy of errors, we come across semantically overlapping error types that are used to evaluate LLM generated explanations. For instance, incorrect reasoning and lack of knowledge are recognized as independent error types in Jin et al. (2024); Hamer et al. (2023); Nievas et al. (2024), even though the latter often leads to the former. In the proposal by Liévin et al. (2024), reasoning errors are separate from reading comprehension, even though the instances of incorrect reading comprehension occur when the model ignores contextual information and “reasons” using its learned knowledge. We therefore identify where errors occur and break them down further by an identifiable source of the error (Figure 2).
Information Isolation Errors
These are errors in information extraction from patient data or trial criteria falls. This includes missed or incorrect NER, measurements (numeric or unit), temporal scopes and negations. LLMs are good at recognizing entities and measurements, but still struggle with negations Nievas et al. (2024).
Reasoning Errors
Any error in the explanations provided by LLMs fall into this category. The most common source is insufficient explicit data. Basically, LLMs fail to draw logical conclusions from given data, while an expert can. This stems from previously unseen data (“lack of knowledge”) or the inability to recall prior information (“lack of implicit reasoning”) or the inability to infer from context (“insufficient data”).
The second source of reasoning errors is lack of self-assessment. The LLM contradicts explicit information in the prompt. The error occurs when knowledge is wrongly recalled or knowledge is correctly recalled, but contextual information is not applied, resulting in wrong reasoning. This is often referred to as “incorrect knowledge”.
The third source are premature decisions made by an LLM when the criteria actually require information that is missing, such as, expert opinion.
Inconsistency Errors
Generating explanations to arrive at final answers unlock LLM’s reasoning capacities Wei et al. (2022). Even so, the explanation and the prediction can be inconsistent. LLMs may predict incorrectly despite correct explanation (reported as explanation-output mismatch in Nievas et al. (2024)). The opposite situation, when a prediction is correct despite an incorrect explanation, is more difficult to report unless explanations for correct predictions are evaluated as well. It may seem as if the LLM is taking a shortcut, hinting towards a bias that the model picked up during training. Both these cases negatively impact transparency and accountability of the matching system.
8 Future Directions
Clinical trial recruitment spans multiple interactive parts. As interest in LLMs orchestrating an end-to-end pipeline and incorporating human interactions is gaining more attention Gao et al. (2024); Qiu et al. (2024), we focus on four promising directions.
Trial Search
Past trials that share a target population are important for designing new trials, recruiting patients, systematic reviews and meta-analyses. This task goes beyond keyword search to match eligibility criteria of two trials. This problem has seen little activity since clustering using lexical features Hao et al. (2014). Analogous systems, such as SemEHR Wu et al. (2018), exist for interactive semantic search over patient EHRs. Sun et al. (2022) combine multiple similarity measures to account for categorical features (study type, phase, and so on) and lexical features (condition name, intervention name, outcome measures, eligibility criteria) to find similar trials. Jullien et al. (2023) deploy LLMs to find trials that match short descriptions using textual entailment and ranked evidences. However, these models falter on inferences that require numerical reasoning and could not surpass a BM25 baseline for ranking evidence.
Interactive Trial Design
Investigators design trials independent of patient data. This creates a semantic gap between eligibility criteria and patient data. LLM agents have the potential of bridging this gap by suggesting data models underlying patient data for structuring eligibility criteria early on. This collaborative idea is not new Luo et al. (2013), however, designing eligibility criteria is a big challenge and has been handled post-hoc by optionally considering patient data models Kang et al. (2017); Sun and Loparo (2019); Liu et al. (2021a); Dasgupta et al. (2020). Small patient pools, which ultimately affect the successful completion of a trial, are often the result of restrictive criteria Clinical Trials Arena (2022). LLMs can improve trial design by identifying restrictive criteria for trial investigators to relax them and create larger patient pools, especially for trials tackling diseases with a high mortality rate Liu et al. (2021b).
Collaborative Trial Planning
Eligibility criteria have a range of difficulty levels. A preemptive task can interact with patient records to estimate the size of the patient pool on the most restrictive criteria, say the presence of a rare genetic disorder. As patient records progressively follow uniform data models, it is more efficient to separate and structure criteria that require tabular operations (e.g., via structured queries) from criteria that require reasoning (temporal, numerical, negation), and to separate criteria that require expert analysis from being wrongly used in semantic searches Mozannar and Sontag (2020). Markey et al. (2025) showed promising results on content relevance and suitability with room for improvement in logical reasoning and provenance when using LLMs to generate trial protocols. Training LLM agents to identify and distribute tasks and then aggregating is a fertile ground for innovation.
Explainable Matches
In the context of LLMs, accurately explaining black-box predictions in human-understandable form becomes very challenging Zhao et al. (2024a). Explainability research in clinical trial matches has been limited to prompting LLMs and manually evaluating chain-of-thought generations post-hoc, which is only one of the many facets of evaluating explainability Nauta et al. (2023); Bodria et al. (2023); Chen and Eickhoff (2024). Moreover, it is unclear how helpful these explanations are, especially, as discussed in Section 7, since they give rise to inconsistency errors and are categorized into ambiguous error labels.
9 Conclusion
The ability of generative LLMs to generalize comes with a trade-off between reasoning and consistency. This work is the first to survey the rapidly evolving role of generative LLMs in clinical trial recruitment. We provide a modern perspective on the challenges in adopting LLMs to clinical research and point out promising directions for future work.
10 Limitations
This survey focuses on the role of advances in NLP in the critical domain of clinical trial recruitment. Given that this is a rapidly evolving field, we have made our best effort to include a comprehensive view of available resources and methods. It is possible that more sophisticated methods using the latest technology already exist (e.g., in the form of proprietary products), but are not yet made public.
Appendix A Trial-Patient Matching Systems
We show all the classical systems in Table 2 and the LLM-based systems in Table LABEL:tab:systems_overview_llm, by their task, method, data source characteristics and limitations.
Appendix B Systematic Reviews versus Surveys
A systematic review is a common practice in the medical research community with standardized reporting guidelines. The Preferred Reporting Items for Systematic Reviews and Meta-Analyses (PRISMA) Vrabel (2015) is the most widely used guideline. Just as finding patients given a trial eligibility criteria, this guideline requires strict inclusion and exclusion criteria to be applied to scientific literature to be included in a systematic review.
In contrast, the guidelines for writing surveys in our community, for instance, as outlined for authors by ACM Computing Surveys , TACL and ARR aim to draw perspectives on an evolving topic of interest.
ACM Computing Surveys (long paper)
A paper that summarizes and organizes recent research results in a novel way that integrates and adds understanding to work in the field. A survey article assumes a general knowledge of the area; it emphasizes the classification of the existing literature, developing a perspective on the area, and evaluating trends.
TACL (excerpt)
They should thus not simply be a descriptive enumeration of the contents of papers, but draw broad themes and (importantly) provide new insights on the topic. These insights should be major contributions of the submission.
ARR (note)
all papers are expected to include reviews of related literature. This category is meant for the papers that go beyond that, e.g. in scope or in establishing new interdisciplinary connections.
Direction Method Patient Source Trial Source Limitations Trial-centric Penberthy et al. (2010) Discrete-data filter and sub-word matching Availability:
Private
Size 282-2,112
Mode: text, tables
Additional annotation: expert review
Availability:
public (1); unknown (4)
Size: 5
Type: Specific
Manually coded eligibility criteria.
String similarity for keyword search.
Yuan et al. (2019) (Criteria2Query)
NL to structured queries
Information extraction, normalization
Mapping to pre-defined cohort templates
No patient data Availability:
Public
Size: 10
Type: Varying
Additional annotation: 125 criteria sentence, 215 entities, 34 relations, 137 negations, 20 attributes
Disease-specific NER.
Query templates cannot deal with missing attributes.
Entity normalization into a small set of 2000 concepts.
Tun et al. (2023)
Rule-based filters
Similarity scores per criteria as features to a classifier
Availability:
Private
Size: 40,000
Mode: text, tables
Additional annotation: 109 patients labels
Availability:
Unknown
Size: 1
Type: Specific
Limited to a single trial
Rules and classifiers are not robust to changes in criteria
Ni et al. (2015)
Discrete-data filter
Index trial and patient bag-of-words vectors
Return top vector matches for a trial
Availability:
Private
Size: 215
Mode: text, tables
Additional annotation: historical match, expert review
Availability:
Public
Size: 55
Type: Specific
12.6% precision
Patient-centric Ni et al. (2015)
Return top vector matches for a patient
(same as previous row) 4% precision
Zhang and Demner-Fushman (2017)
Bag-of-words feature vector
SVM classifier
No patient data Availability:
Public
Size: 2461
Type: 891 Specific; 1570 Varying
Additional annotation: Trial labels
Cohort-specific model
No real patient data considered
Closest to keyword search (trials to cohort)
Public source for clinical trials: https://clinicalTrials.gov
Table 2: Overview of classical systems covered in this survey: direction, methods, data sources and limitations.
Task Method Patient Source Trial Source Limitations
Trial-level prediction Wong et al. (2023)
LLM prompted with a template to generate structured forms of eligibility criteria
Availability: Private
Size: Unknown
Mode: Structured
Additional Annotation: 523 trial-patient historical labels; 68,485 trial-patient new labels
Availability: Public∗∗
Size: 53
Type: Specific
Clinical Trial eligibility limited to the first 40 lines.
Trial-level performance is very low (F1 score range:[29.6-48])
Yuan et al. (2023) Prompt LLMs to reformulate eligibility criteria Availability: Private
Size: 825
Mode: Longitudinal text records
Availability: Public∗∗
Size: 6
Type: Specific
Additional Annotation: 100,000 criterion-patient labels
Criterion-level performance metrics 10 points higher than trial-patient.
High variance between different trials: F1 range: 0.39 - 0.96.
Criterion-level prediction Hamer et al. (2023)
1-shot prompt to LLM for criteria level prediction with explanation
Availability: Private (synthetic patient profile)
Size: 10
Mode: Short text
Type: Specific
Availability: Mixed (clinicalTrials.gov, EudraCT)
Size: 146 clinical trials
Type: Specific
Additional Annotation: Expert review
Majority of criterion-level errors due to incorrect reasoning (91%)
Unlu et al. (2024)
RAG-based
Availability: Private
Size: 2,276
Mode: Tables; text
Availability: Public∗∗
Size: 1
Type: Specific
RAG pipeline not published. Evaluated on a single trial.
Wornow et al. (2025)
Vector database of patient notes
Prompt (criteria, top-k patient notes) pair to LLM
Availability: Public (2018 N2C2: Cohort Selection) Evaluated on one trial
Doesn’t consider exclusion criteria
67% incorrect decision despite correct reasoning.
Trial ranking Pradeep et al. (2022)
Synthesis queries for initial trial retrieval.
Fine-tune T5 to generate relevance label based on patient description and trial data
Availability: Public (TREC CT 2021) Zero-shot relevance ranking is only slightly better than BM25. Fine-tuned models need re-training on new data or format.
Zhuang et al. (2023)
Hybrid sparse-dense retriever for top-1000.
Cross-encoder re-ranker
GPT4 for top-20 re-rank
Availability: Public (TREC CT 2023) High NDCG, 0.51 precision, 0.38 recall (very low)
Validation on textual patient note does not transfer well to testing on structured patient note
Rybinski et al. (2024)
Multi-stage retriever using LLMs
Availability: Public (TREC CT 2021, 2022, 2023) High latency of GPT4 vs a small performance boost
Jin et al. (2024)
Hybrid trial filtering with BM25 and MedCPT Jin et al. (2023)
Criterion-level LLM prediction aggregated to trial scores
Availability: Public (Koopman and Zuccon (2016), TREC CT 2021 & 2022)
Size: 58; 75; 50
Mode: Short text
Additional annotations: 3,621; 26,149; 26,581 trial-patient matches
Availability: Public∗∗
Size: 375,580
Additional annotation: 1,015 criterion-patient labels
User study sample size is small
Does not handle longitudinal patient data
LLM aggregation requires LLM to perform mathematical reasoning and computations
Nievas et al. (2024)
Extends Jin et al. (2024) to open-source LLMs
(same as Jin et al. (2024))
Additional annotation: Patient sentence supporting eligibility label Additional annotation: 500 criteria labels on: eligibility and difficulty.
Significantly high fine-tuning costs
EudraCT (European Union Drug Regulating Authorities Clinical Trials) is the European clinical trials database EudraCT .
∗∗ Public source for clinical trials: clinicalTrials.gov.
Table 3: Overview of LLM-based systems covered in this survey: tasks, methods, data sources used and limitations.

### v2

@@
-Cohort Discovery: A Survey on LLM-Assisted Clinical Trial Recruitment
+A Survey on LLM-Assisted Clinical Trial Recruitment
 Abstract
-Recent advances in LLMs have greatly improved general-domain NLP tasks.
-Yet, their adoption in critical domains, such as clinical trial recruitment, remains limited.
-As trials are designed in natural language and patient data is represented as both structured and unstructured text, the task of matching trials and patients benefits from knowledge aggregation and reasoning abilities of LLMs.
-Classical approaches are trial-specific and LLMs with their ability to consolidate distributed knowledge hold the potential to build a more general solution.
-Yet recent applications of LLM-assisted methods rely on proprietary models and weak evaluation benchmarks.
-In this survey, we are the first to analyze the task of trial-patient matching and contextualize emerging LLM-based approaches in clinical trial recruitment.
-We critically examine existing benchmarks, approaches and evaluation frameworks, the challenges to adopting LLM technologies in clinical research and exciting future directions.
-Cohort Discovery: A Survey on LLM-Assisted Clinical Trial Recruitment
+Clinical trials are designed in natural language and the task of matching them to patients, represented via both structured and unstructured textual data, benefits from knowledge aggregation and reasoning abilities of LLMs.
+LLMs with their ability to consolidate distributed knowledge hold the potential to build a more general solution than classical approaches that employ trial-specific heuristics.
+Yet, adoption of LLMs in critical domains, such as clinical research, comes with many challenges, such as, the size of the benchmarks, the dimensions of evaluation and data sensitivity.
+In this survey, we contextualize emerging LLM-based approaches in clinical trial recruitment.
+We examine the main components of the clinical trial recruitment process, discuss existing challenges in adopting LLM technologies in clinical research and exciting future directions.
+A Survey on LLM-Assisted Clinical Trial Recruitment
 Shrestha Ghosh1    Moritz Schneider2    Carina Reinicke2    Carsten Eickhoff1 1University of Tübingen, Germany 2Boehringer Ingelheim, Germany 1{first.last}@uni-tuebingen.de   2{first.last}@boehringer-ingelheim.com
@@
 (2015); Pharmaceutical Technology (2019).
-While the trend has steadily declined over the past decade, with the intensive use of technology-aided solutions, efficient patient recruitment remains the most crucial bottleneck in clinical trial research Clinical Trials Arena (2022), thereby hampering the rate at which potentially life-saving therapies or new vaccines can be introduced.
-Figure 1: Steps in a typical patient recruitment process vs. our proposed flow with expert interaction.
-Clinical trials have a dual nature, consisting of both trial-agnostic and trial-specific requirements, making it challenging to design generalized approaches.
-This has traditionally resulted in highly specialized matching processes that work for a particular trial or patient population.
-The resulting patient recruitment process is linear, as illustrated (in orange) in Figure 1, leaving little room for expert intervention and feedback.
-It starts with extracting information from a trial description and available patient records, then matching explicit trial eligibility criteria and patient data to determine patient eligibility.
-Finally, experts review the selected patients before contacting them.
-An expert reviews several hundred patients Ni et al.
-(2015) and can end up spending hours on one patient, incurring significant costs Penberthy et al.
-(2012).
-In their study on five trials, Penberthy et al.
-(2010) showed that automated matching, via table queries and string-matching algorithms, saved between 165 hours to 1,329 hours of reviewing time when compared to manual evaluation.
-As more patient data becomes digitally available, research into applying machine intelligence and developing explainable systems for clinical researchers to correctly interpret model predictions has gained traction Murdoch and Detsky (2013); Payrovnaziri et al.
+While the trend has steadily declined over the past decade, with the intensive use of technology-aided solutions, efficient patient recruitment remains the most crucial bottleneck in clinical trial research Clinical Trials Arena (2022).
+As electronic health records (EHRs) of patients become more accessible, clinical researchers adopt machine intelligence and develop explainable systems to correctly interpret model predictions Murdoch and Detsky (2013); Payrovnaziri et al.
 (2020); von Itzstein et al.
 (2021).
-Generative LLMs can serve as knowledge aggregators and streamline clinical trial design and trial-patient matching (in purple in Figure 1).
-However, the clinical trial recruitment process has seen relatively low adoption of transformer language models He et al.
+There are methods leveraging LLMs in clinical trial research for cohort retrieval and modelling Fang et al.
+(2022); Tian et al.
+(2023); Park et al.
+(2024); Liu et al.
+(2025a); Wang et al.
+(2025), trial design Reinisch et al.
+(2024); Curran et al.
+(2024); Bornet et al.
+(2025); Neehal et al.
+(2025), trial search White et al.
+(2023); Rybinski et al.
+(2020), trial matching Jin et al.
+(2024); Nievas et al.
+(2024); Wornow et al.
+(2025), predicting trial outcomes and duration Reinisch et al.
+(2024); Yue et al.
+(2024a, b); Liu et al.
+(2025c), assessing risk of bias Lai et al.
+(2024); Ji et al.
+(2025), and extracting clinical trial results Lee et al.
+(2024), even as the community is coming up with recommended practices for responsible use of AI in the entire drug development process Geraci et al.
+(2025).
+Figure 1: Components in a patient recruitment process: conventional linear flow (in orange) vs. our proposed LLM-assisted interactive flow (in purple).
+Figure 1 shows the components in clinical trial recruitment, namely, data sourcing, information extraction, matching, and evaluation.
+An expert reviews several hundred patients per trial and can end up spending hours on one patient, hence incurring significant costs Penberthy et al.
+(2012); Ni et al.
+(2015).
+Even simple automation using table queries and lexical searches saves between 165 hours to 1,329 hours of reviewing time when compared to manual evaluation Penberthy et al.
+(2010).
+In the past, the patient recruitment process has seen relatively low adoption of the pretrained language models He et al.
 (2020); Harrer (2023); Lu et al.
 (2024).
-Instruction fine-tuned LLMs for generating explainable trial-patient matches has received attention only since the past one year Jin et al.
+Generative LLMs serve as knowledge aggregators, and through their reasoning and instruction-following capabilities, they have revived research in the task of trial and patient matching Jin et al.
 (2024); Nievas et al.
@@
 (2025).
-Given the sensitive nature of clinical trials and the rapidly evolving landscape of LLM technology, there is no prior work surveying this area.
-We fill this gap by critically examining the trial-patient matching approaches, the method used, the benchmarks available and the evaluation strategies.
-The existing systematic reviews of NLP methods and their performance Kim and Quintana (2022); Idnay et al.
-(2022); Kantor and Morzy (2024) are desiged to follow strict inclusion/exclusion criteria and search strategies Khan et al.
-(2003); Moher et al.
-(2009).
-This practice, essential in medical research for measuring the association between intervention and outcomes, misses the larger picture of contextualizing LLMs in clinical trial recruitment settings.
-In contrast, this work follows the perspective of surveys in computer science (see Appendix B) to provide a more comprehensive view of the role of LLMs in clinical trial recruitment.
+Difference to Prior Work
+Despite the rapidly evolving landscape of LLM technology, there is no prior work surveying this area.
+Gueguen et al.
+(2025) evaluate public trial matching tools and Layne et al.
+(2025) compare the efficacy of open and proprietary LLM-assisted trial-patient matching in oncology.
+Systematic reviews in this topic, bound by strict selection criteria and highly specific research question, do not to capture the broad perspectives Kim and Quintana (2022); Idnay et al.
+(2022); Kantor and Morzy (2024b); Chen et al.
+(2025) (Appendix A differentiates between systematic reviews and surveys).
+Although it provides an overview of LLM approaches used in clinical trial matching, Chen et al.
+(2025) only briefly discusses associated challenges and future work.
 Our Contribution
-Cohort discovery has been a long-standing problem and LLM-assisted methods to improve trial-patient matching has only recently gained momentum.
-In this work, we formalize the problem of trial-patient matching, and differentiate between two major approaches (trial-centric and patient-centric), the associated tasks (classification vs.
-ranking) and their effect on the choice of benchmarks (longitudinal vs.
+We examine the main components in a trial recruitment process, presented in Figure 1.
+We formalize the problem of trial-patient matching.
+We analyze existing approaches via the tasks (classification vs.
+ranking), the directionality (trial-centric and patient-centric) and the choice of benchmarks (longitudinal vs.
 short patient descriptions, single vs.
 multi-trial).
-We identify the important bottlenecks in data availability and task formalization.
-We introduce a taxonomy of errors in LLM-generations for trial-patient matching.
-Finally, we provide actionable steps for the community on four promising directions.
+We present a taxonomy of errors for consistent evaluation of LLM-generated responses.
+We discuss the critical challenges associated with using LLMs in trial recruitment research.
+Finally, we present actionable steps towards an interactive patient recruitment (illustrated in purple in Figure 1).
 2 Background
 Clinical Trial Recruitment
-Also known as patient recruitment/enrollment/(pre-)screening, it is the process of finding relevant patients (or cohort) for a clinical trial.
-The steps in patient recruitment, as shown in Figure 1, include data sourcing, information extraction, matching, and expert review.
-Eligibility Criteria
-They are explicit conditions for (inclusion) or against (exclusion) enrollment, usually expressed in natural language and designed by trial investigators to select cohorts.
-These conditions have varying difficulty levels, ranging from simple tabular operations (age or gender requirements) to semantic matching (medical condition or medications) and inference (reasoning with negation, patient history).
-Patient Records
-They refer to any form of data, in a combination of structured and unstructured formats, used to describe a patient.
-When digitized, it is often called an Electronic Health Record (EHR).
-This data includes the patient’s demographic information (age, gender, ethnicity), clinical notes, medications, billing history, and test results, that are collected over a period of years.
+Also known as patient recruitment/enrollment/(pre-)screening, it is the process of matching patients (or a cohort) to a clinical trial via its eligibility criteria.
+Clinical trials have a dual nature, consisting of both universal and trial-specific requirements, making it challenging to design generalized approaches Idnay et al.
+(2023).
+This has traditionally resulted in linear trial-centric matching processes (illustrated in orange in Figure 1) with limited scope for interaction and feedback.
+Such systems generally have an initial data filter on the structured EHRs followed by keyword matches and concept identification Penberthy et al.
+(2010); Tun et al.
+(2023) or cohort-specific classifiers Zhang and Demner-Fushman (2017).
+Ni et al.
+(2015) represented both trials and patients as feature vectors, thus supporting retrievals for both trial and patient.
 Biomedical NLP and LLMs
-Transformer language models have greatly improved information retrieval and extraction tasks.
-There exist specialized resources, such as, PubmedQA Jin et al.
-(2019), MIMIC-III & IV Johnson et al.
-(2016, 2020) and BioRED Luo et al.
-(2022), to train encoder language models Radford (2018); Devlin et al.
-(2019); Raffel et al.
-(2020) on medical domain (e.g., BioBERT Lee et al.
+The biomedical NLP landscape is shifting from specialized pretrained language models, such as, BioBERT Lee et al.
+(2020), BioLM Lewis et al.
 (2020), PubmedBERT Gu et al.
-(2021), Criteria2Query2.0 Fang et al.
+(2021), BioGPT Luo et al.
 (2022), MedCPT Jin et al.
-(2023)).
-Yet, Dorfner et al.
-(2024) show that fine-tuning generative LLMs on biomedical domain has limited relative performance gain.
+(2023), among many others Wang et al.
+(2023), towards instruction-following and chat-enabled LLMs (commonly termed as generative AI), used as is Nori et al.
+(2023); Kung et al.
+(2023) or fine-tuned for domain alignment, such as, Med-PaLM Singhal et al.
+(2023), Med-Alpaca Han et al.
+(2023) and LLaVA-Med Li et al.
+(2023).
+We point our reader to Thirunavukarasu et al.
+(2023); Liu et al.
+(2025b) for further reading.
+The Journal of American Medical Informatics Association (JAMIA), which published 41 articles on biomedical health and LLMs in a focus issue, also observed this shift towards generative AI Lu et al.
+(2024).
+Among these, the OpenAI Achiam et al.
+(2023) family of proprietary models (GPT3.5 and GPT4 being the most common) is used far more often than open-sourced models Touvron et al.
+(2023); Bai et al.
+(2023); Jiang et al.
+(2023) models.
+Further, fine-tuning generative LLMs on the biomedical domain offers limited performance gain Dorfner et al.
+(2024).
 In fact, Alber et al.
-(2025) show that medical misinformation encountered during the fine-tuning affect downstream tasks.
-The Journal of American Medical Informatics Association (JAMIA), which published 41 articles on biomedical health and LLMs in a focus issue, observed a shift from domain-specific small models towards larger general-purpose LLMs Lu et al.
+(2025) show that models are more prone to propagate medical misinformation encountered during fine-tuning despite performing well on benchmarks.
+Finally, Bedi et al.
+(2025) report that real patient data is used in less than 5% of the 519 biomedical studies using LLMs, and fairness, bias, uncertainty and deployment considerations are rarely assessed.
+State of Adoption of NLP Advancements
+Kantor and Morzy (2024b)’s study on AI adoption rates in eligibility criteria parsing, report very low adoption rates of generative AI, with BERT family of models being most popular and generative models coming into the picture only since 2024.
+A systematic review of the role of NLP systems in patient recruitment identified only 11 studies Idnay et al.
 (2022).
-Even among LLMs, the OpenAI Achiam et al.
-(2023) family of proprietary models (GPT3.5 and GPT4 being the most common) is used far more often than open-sourced LLaMA Touvron et al.
-(2023) or Mistral Jiang et al.
-(2023) models.
-Finally, Bedi et al.
-(2025) report that of more than 519 biomedical studies using LLMs, less than 5% used real patient data, tasks like summarization are understudied and fairness, bias, uncertainty and deployment considerations are rarely measured.
-State of Adoption of NLP Advancements
-Historically, patient recruitment has been a manual process with minimal automation via pre-defined database queries Thadani et al.
-(2009), and, more recently, automatic translations of free text into database queries Yuan et al.
-(2019); Fang et al.
-(2022); Park et al.
-(2024).
-A recent review reported low adoption rates of AI, particularly NLP, in clinical trials Kantor and Morzy (2024).
-Another review on understanding the role of NLP systems in patient recruitment found just 11 studies that evaluated NLP systems in clinical settings Idnay et al.
-(2022).
-Heterogeneous outcomes, diverse results and a dependence on small retrospective data drive the gap in NLP research and their adoption in real-world settings Idnay et al.
-(2024).
+Heterogeneous outcomes, diverse results, a dependence on small retrospective data and a lack of common standardized benchmarks drive the gap in NLP research and their adoption in real-world settings Idnay et al.
+(2024); Kantor and Morzy (2024b).
 A study by Idnay et al.
@@
 Interestingly, when Corbaux et al.
-(2024) categorize trial matching tools for oncological trials as manual, semi-automatic and automatic, automatic methods still fall in the research and development phase, yet to be commercially available.
-Review Methodology
-We searched for relavant papers since 2019, using the keywords “clinical trial”, “cohort discovery”, “patient recruitment”, “trial recruitment”, “trial matching” in conjunction with “llm”, “language model”, “gpt”.
-We included method and dataset papers using LLMs and excluded papers that used LLMs in biomedical NLP (broader scope) or other health care applications (not clinical trials).
-Starting with the citations in the LLM papers we tracked the pre-LLM literature by recursively tracing their citations in newer publications.
-Since the topic “clinical trial recruitment” is vast, this approach efficiently captured the evolution of the predictive methods in this field.
-3 Trial-Patient Matching
-Trial-patient matching assigns a trial-patient pair one of the three labels of eligible, ineligible and irrelevant based on the patient record and eligibility criteria.
-Importantly, a trial-centric approach, taken by a trial investigator, matches patients from a hospital database to a specific trial.
-Alternatively, a patient-centric approach, taken by a patient or their healthcare provider, matches relevant trials from a trial registry to a patient.
-Given sets of inclusion and exclusion criteria ($C_{inc},C_{exc}$) from a trial and a set of patient data, $P$, we formalize the trial-patient matching problem, $M$, to predict one of the labels Ineligible (Inel.), Irrelevant (Irr.) or Eligible (Eli.) by aggregating criterion-level binary matches $M^{\prime}(c,P)$.
-$\!M(C_{inc},C_{exc},P)\!=\!\left\{\!\begin{array}[]{ll}\!\textit{Inel.},\!&\!\exists c\in C_{exc},M^{\prime}(c,P)\\ \!\textit{Irr.},\!&\!\exists c\in C_{inc},\neg M^{\prime}(c,P)\\ \!\textit{Eli.},\!&\!\neg(\textit{Inel.}\lor\textit{Irr.})\end{array}\!\right.$
-This induces a priority, such that, a patient satisfying any exclusion criteria becomes ineligible, regardless of inclusion criteria matches.
-In Section 6, we elaborate on the lack of a formal task definition and its effects on trial-centric matching.
-Task Benchmark Size Eval. Metrics Best Score Cohort
+(2024) categorize tools for oncological trial matching, they indicate that the automatic methods still fall in the research and development phase, yet to be commercially available.
+3 Methodology
+We queried Google Scholar with the keywords “clinical trial”, “cohort discovery”, “patient recruitment”, “trial recruitment”, “trial matching” in conjunction with “llm”, “language model”, “gpt” and inspected the first ten result pages, between the years 2019 and 2025 (both inclusive).
+We also used Jin et al.
+(2024) and Wornow et al.
+(2025) as seeds and recursively traced their citations to efficiently captured the evolution of the predictive methods in this field.
+We organized papers that described LLM-based approaches and corpora for clinical trial recruitment into data sourcing (Section 2), information extraction and parsing (Section 5), trial-patient matching (Section 6) and evaluation (Section 7).
+Via our discussion on critical limitations in Section 8 and on promising directions towards an interactive patient recruitment in Section 9, we provide further contextualize LLMs in healthcare settings and other stages of the drug development process that directly effect on patient recruitment.
+Task Benchmark Size Eval. Metrics Best Score #P LLM Usage (#P) Cohort
 selection
@@
 P, R, F1
-0.91 micro F1
+0.91 micro F1 47 (not applicable)
 Trial
@@
 #relevant matches = 685
-#relevance judgements = 4000
+#relevance judgments = 4000
 MRR,
@@
 0.3 MRR
-$<$ 0.2 P@5
-TREC CT Track 2021
+<< 0.2 P@5
+- $\dagger$ (not applicable)
+TREC CT 2021
 Soboroff (2021)
@@
 #relevant matches = 5,570
-#relevance judgements = 35,832
+#relevance judgments = 35,832
 NDCG@10,
@@
 1.0 MRR
-TREC CT Track 2022
+26 BERT-based keyword extraction (9) / query summarization (1),
+Transformer-based rankers (10)
+TREC CT 2022
 Roberts et al. (2022)
@@
 #relevant matches = 3,949
-#relevance judgements = 35,394
+#relevance judgments = 35,394
 NDCG@10,
@@
 0.72 MRR
-TREC CT Track 2023
+12 Query reformulation using BERT / sequence-to-sequence models (3),
+Transformer-based rankers (3)
+TREC CT 2023
 Rybinski et al. (2024)∗∗
@@
 #relevant matches = 11,963
-#relevance judgements = 34,931
+#relevance judgments = 34,931
 NDCG@10,
@@
 0.78 MRR
-Best scores are aggregated from the Appendix of the proceedings of TREC 2021. ∗∗Borrowed from Rybinski et al. (2024) as TREC CT 2023 does not have a published track overview.
-Table 1: Overview of public evaluation benchmarks for trial-patient matches covered in this survey.
-4 Benchmarks and Corpora
-We analyze five public benchmarks in context of the corpora and their evaluation metrics.
-  • 2018 N2C2 Cohort Selection for Clinical Trials Stubbs et al. (2019). [^1] Currently unavailable as of 2024 Nov 6.
-  • Trial ranking Koopman and Zuccon (2016).
+11 Query reformulation using LLMs (5),
+Transformer-based rankers (6),
+LLM prompt-based relevance prediction (4)
+Best scores are aggregated from the Appendix of the proceedings of TREC 2021.
+No participants, since this was not a challenge.
+∗∗Borrowed from Rybinski et al.
+(2024) as TREC CT 2023 does not have a published track overview.
+Table 1: Overview of the public trial-patient matching benchmarks.
+#P is the number of participating teams.
+LLM usage (#P) tracks the number of participants using LLMs that we could verify from the proceedings.
+4 Data Sourcing: Public Benchmarks
+We analyze five trial-patient matching benchmarks.
+  • 2018 N2C2 Cohort Selection Stubbs et al. (2019) for criterion-level eligibility prediction. [^1] Currently unavailable as of 2024 Nov 6.
+  • Koopman and Zuccon (2016) for ranking trials.
   • Text REtrieval Conference Clinical Trial (TREC CT) [^2] Available at https://www.trec-cds.org/. tracks 2021, 2022, 2023 for ranking trials.
@@
 4.1 Trial Data
-ClinicalTrials.gov is one of the largest online database of clinical studies, including clinical trials and other observational studies.
-As of November 2024, it lists more than 390,000 clinical trials, thousands of which are active.
-These studies are submitted by investigators from over 200 countries.
-The euclinicaltrials.eu is another online registry for trials conducted in the European Union.
-It comprises more than 50,000 trials, with more than 7000 being active.
+The TREC CT benchmarks and Koopman and Zuccon (2016) source the ClinicalTrials.gov registry.
+Meanwhile, the N2C2 benchmark focuses on a single trial.
+ClinicalTrials.gov is one of the largest online databases of clinical studies submitted by investigators from over 200 countries.
+As of June 2025, it lists more than 400,000 clinical trials, thousands of which are active.
+The euclinicaltrials.eu is another online registry comprising over 50,000 trials from the European Union of which, 7000 are active.
 Every trial comprises a study description, eligibility criteria and study plan among other details.
-The ClinicalTrials.gov database forms the primary source of trials in many academic studies.
-As we will see in Section 5, earlier trial-specific approaches manually separated discrete data from free-text search.
-However, this separation is difficult to scale, and newer approaches focus on semantic matching.
-The N2C2 cohort selection benchmark is based on a single trial.
-Hence, systems specifically model the trial criteria.
-The TREC CT benchmarks source the ClinicalTrials.gov database for the task of trial ranking.
-However, additional annotation of the eligibility criteria denoting their type or complexity is unavailable.
 4.2 Patient Data
-Since patient data contains sensitive information, additional measures are required to remove personal information in order to release them as public repositories.
+The N2C2 benchmark contains 288 de-identified longitudinal records of patients.
+The trial ranking benchmarks use up to 75 synthetic patient profiles, comprising keywords defining cohorts Koopman and Zuccon (2016), short textual patient descriptions (TREC CT 2021, 2022) and questionnaire templates (TREC CT 2023).
 The MIMIC database Johnson et al.
-(2016, 2020) is the largest public database of patient information.
-The latest version, MIMIC-IV, contains rich data on hospital admissions, medications, demographic data and notes of 299,712 de-identified patients.
-This dataset is not used in any of the five benchmarks.
-One reason might be a significant challenge in obtaining confirmed trial-patient match labels, which is easily available retrospectively from existing hospital records.
-The N2C2 cohort selection benchmark Stubbs et al.
-(2019), contains 288 de-identified longitudinal records of patients.
-The TREC CT benchmarks use synthetic patient profiles, comprising short textual patient descriptions and keywords defining cohorts.
-4.3 Evaluation Metrics
-Standard IR metrics (precision, recall, F1) and their derivatives (precision at rank (P@k), normalized discounted cumulative gains (NDCG@k), R-precision, mean reciprocal rank (MRR) and adaptive precision Moffat et al.
-(2015)) are the most popular evaluation metrics.
-Accuracy, specificity, and workload are also prevelant in the medical community.
-Notably, researchers in the medical domain have stressed the lack of definition and consistency of ground truth data to evaluate NLP tools in clinical trial matching Kim and Quintana (2022).
-The trial-patient ground-truth labels are obtained manually either from medical experts Koopman and Zuccon (2016); Soboroff (2021); Roberts et al.
-(2022) or retrospectively from past trials Stubbs et al.
+(2016, 2020) is the largest anonymized public database of structured and unstructured data of 299,712 patients.
+While this dataset is popular for training biomedical LLMs (as mentioned in Section 2), it is not used in any of the five benchmarks, possibly due to significant challenges in obtaining eligibility labels at this scale.
+4.3 Annotated Labels
+Two medical experts annotated 3,744 criterion-level labels in the N2C2 benchmark.
+The TREC CT annotations were created by pooling the top-k results from all participating teams.
+Medical experts manually annotated this pool of trial-patient matches Koopman and Zuccon (2016); Soboroff (2021); Roberts et al.
 (2022).
-The TREC CT annotations were created from the top-k results pooled among all participating teams, resulting in tens of thousands of relevance judgments.
 Out of a total of 35,394 pooled trial-patient matches in the 2022 edition, 11% were judged as eligible, 9% as ineligible and 80% as not relevant with an average of 700 trials judged per patient Roberts et al.
 (2022).
-In the 2023 edition, the relevance judgments were more balanced between the three labels Rybinski et al.
+The relevance judgments were more balanced between the three labels Rybinski et al.
+(2024) in TREC CT 2023.
+All benchmarks depend on manual annotation from experts, which is time-consuming and challenging to scale Kim and Quintana (2022).
+5 Information Extraction and Parsing
+Extraction of medical entities evolved from a combination of rule-based heuristics and feature-based supervised sequence labelling models Kang et al.
+(2017); Yuan et al.
+(2019) via embedding-based neural models Khan et al.
+(2019); Tseo et al.
+(2020) to transformer-based pretrained models and LLMs Liu et al.
+(2021c); Zeng et al.
+(2020); Tian et al.
+(2021); Li et al.
+(2022); Murcia et al.
+(2024); Kantor and Morzy (2024a).
+Datta et al.
+(2024) use disease-specific prompting to extract structured information about entities and its attributes from criteria text.
+Gao et al.
+(2020); Zhang et al.
+(2020); Theodorou et al.
+(2023) use BERT-based embeddings to encode eligibility criteria and patient data.
+Patient data converted to search queries, via reformulation and expansion using LLMs, are particularly effective in trial retrieval Peikos et al.
+(2023); Rybinski et al.
+(2024); Peikos et al.
+(2024); Jin et al.
 (2024).
-Measurement of human preference of LLM generations, bias in model predictions and the informativeness of LLM explanations ranges from limited to non-existent in the current works.
-5 Trial-Patient Matching Methods
-We broadly categorize systems into classical and LLM-based approaches and further group them by their direction of approach, trial/patient-centric.
-The LLM-based approaches focus on tasks, such as, criterion/trial-level predictions and relevance ranking, which in principle is direction-agnostic.
-Appendix A lists the works cited in this paper.
-5.1 Classical Approaches
-Trial-centric Matches
-Majority of classical models use this approach build trial or cohort specific models.
-They follow a two-step method, starting with manually-designed rule-based discrete data filters, followed by sub-word matching Penberthy et al.
-(2010), vector matching Ni et al.
-(2015) or machine-learned classifiers Tun et al.
-(2023) for matching eligibility criteria or cohort keywords.
-Ni et al.
-(2015) showed a retrospective workload reduction to 50 patients down from 163 per trial.
-IBM Watson for clinical trial matching (IBM-CTM) is a proprietary commercial tool for trial-patient matching, which uses neural NLP techniques to match eligibility criteria to patients Helgeson et al.
+Yuan et al.
+(2019); Tian et al.
+(2023); Park et al.
+(2024); Mugambi et al.
+(2024); Ziletti and D’Ambrosi (2025) use semantic parsing to translate eligibility criteria into logical forms ready for querying patient databases.
+6 Trial-Patient Matching
+6.1 Formalization
+Given sets of inclusion and exclusion criteria ($C_{inc},C_{exc}$) from a trial and a set of patient data, $P$, we formalize the trial-patient matching problem, $M$, to predict one of the labels Ineligible (Inel.), Irrelevant (Irr.) or Eligible (Eli.) by aggregating criterion-level binary matches $M^{\prime}(c,P)$.
+$M(C_{inc},C_{exc},P)\!=\!\left\{\!\begin{array}[]{ll}\!\textit{Inel.},\!&\!\exists c\in C_{exc},M^{\prime}(c,P)\\ \!\textit{Irr.},\!&\!\exists c\in C_{inc},\neg M^{\prime}(c,P)\\ \!\textit{Eli.},\!&\!\neg(\textit{Inel.}\lor\textit{Irr.})\end{array}\!\right$.
+This induces a priority, such that, a patient satisfying any exclusion criteria becomes ineligible, regardless of inclusion criteria matches.
+Section 6.2 explores the task at varying levels of granularity, starting from criterion-level via trial-level predictions to trial ranking.
+In Section 8, we elaborate on the importance of formalization and its effects on trial-level aggregation.
+Despite the task being direction-agnostic, there exist two directional approaches to tackle the matching problem, mainly due to data availability.
+First is the trial-centric approach, taken by a trial investigator, that matches longitudinal patient records to a specific trial.
+The 2018 N2C2 cohort selection is a trial-centric benchmark with criterion-level predictions.
+Alternatively, a patient-centric approach, taken by a patient or their healthcare provider, matches relevant trials from a trial registry to a short description of the patient.
+Koopman and Zuccon (2016) and the TREC CT 2021, 2022, and 2023 are patient-centric benchmarks for ranking trials.
+Classical LLM-based !
+Direction-specific approaches, applicable to a set of trials or a cohort * Direction-agnostic criteria- & trial-level prediction !
+Trial-specific heuristics: filters and features * Generalizable across trials !
+Evaluated on private patient data * Public benchmarks more common
+Table 2: Differences between classical and LLM-based trial matching approaches.
+6.2 LLM-based Approaches
+Unlike classical approaches Penberthy et al.
+(2010); Zhang and Demner-Fushman (2017); Yuan et al.
+(2019); Tun et al.
+(2023), LLM-based approaches do not rely on trial-specific heuristics.
+Table 2 lists the primary differences between the classical and LLM-based approaches (see Appendix C for a full comparison).
+We group the LLM-based approaches by the granularity of the matches, starting with criterion-level prediction via trial-level prediction to trial ranking.
+While criterion- and trial-level predictions are direction-agnostic, trial ranking is a patient-centric task.
+With enough patients, we could evaluate patient ranking, similar to the bidirectional implementation in Ni et al.
+(2015)’s work, though this has not yet been addressed.
+Criterion-Level Prediction
+Here, methods utilize the reasoning capability of LLMs to obtain decision rationale and other context data in addition to the eligibility label.
+Hamer et al.
+(2023) use 1-shot prompt, where given the patient profile and the eligibility criteria, the LLM first labels each criterion as being applicable to the patient or not, followed by a list of rationales and finally, the eligibility labels for the applicable criteria.
+Unlu et al.
+(2024); Beattie et al.
+(2024) and Wornow et al.
+(2025) chunk longitudinal patient data and store them as vector embeddings.
+They prompt LLMs with criteria and relevant patient data chunks under zero-shot setting, with Beattie et al.
+(2024) providing expert criterion-level strategy in the prompts.
+Unlu et al.
+(2024) generate only the decision label from GPT4 and Beattie et al.
+(2024) and Wornow et al.
+(2025) generate criterion-level JSON objects, comprising the criteria label, the eligibility label and a rationale among other context.
+Both work with proprietary OpenAI models (GPT-3.5 and GPT-4) and Wornow et al.
+(2025) additionally used open-sourced Llama Jiang et al.
+(2023) and Mixtral models Jiang et al.
 (2024).
-While one study on 3 trials measured a reduction in workload from 110 minutes to 86 minutes for screening 90 patients Beck et al.
-(2017), another study on 10 clinical trials and 102 patients evaluated the agreement in matches between IBM-CTM and expert annotation to be above 80% Alexander et al.
-(2020).
-Another line of work that focuses on defining a cohort by translating natural language criteria into formal query language for instantly querying patient databases Yuan et al.
-(2019).
-Patient-centric Matches
-Ni et al.
-(2015) create bag-of-words vectors to retrieve top trial matches for a patient and show a retrospective workload reduction (4 vs.
-42 trials per patient) and 33% precision on a subset of 127 patients.
-The workload increases to 20 trials per patient and the precision drops to less than 4% when generalized to 215 patients.
-Zhang and Demner-Fushman (2017) train an SVM classifier on the top unigram and bigram features to predict if a trial is eligible for a cohort (e.g., HIV patients).
-Additional features derived from NER and lemmatization yielded small gains in performance.
-5.2 LLM-based Approaches
-These approaches do not rely on trial-specific heuristics.
-Yet, as systems are evaluated on a few disease-specific trials, the largest being 146 trials, covering 10 cancer types, evaluated by Hamer et al.
-(2023), the generalizability of LLMs across diseases is unclear.
-Moreover, when compared, LLMs are less effective in making trial-level predictions (low F1 scores) than criterion-level ones.
+Ferber et al.
+(2024) prompt GPT-4o to generate criterion-level boolean predictions to be reviewed by experts.
 Trial-Level Prediction
 Wong et al.
-(2023) used GPT4 to structure eligibility criteria into logical expressions.
-Despite lagging behind expert systems, which use heavy manual intervention, they demonstrate the readiness of general-purpose LLMs for cohort-specific NER (oncology in this case) and logical structuring.
-Zero-shot and few-shot LLM prompts outperformed the state-of-the-art biomedical NER, SciSpaCy Neumann et al.
+(2023) used GPT3.5 and GPT4 to extract and structure eligibility criteria into logical expression to be matched locally with structured patient information.
+Yuan et al.
+(2023) use LLMs to augment eligibility criteria and pass the BERT-based embeddings of these criteria and patient data through a fully connected classification layer to predict patient-criterion eligibility.
+The trial-level prediction is then computed by asserting Equation 6.1.
+Trial Ranking
+Here, methods formulate effective queries, retrieve trials and re-rank them.
+Query processing techniques involve generating sentence queries from patient descriptions using a fine-tuned T5 model Pradeep et al.
+(2022) or zero-shot LLM prompts Saeidi et al.
+(2023); Kusa et al.
+(2023b), generating patient descriptions from trial data via 1-shot LLM prompts Zhuang et al.
+(2023), generating no-SQL queries via LLMs Ferber et al.
+(2024), using LLMs to reformulate and expand queries Rybinski et al.
+(2024); Peikos et al.
+(2024); Datta et al.
+(2025), and using LLMs to extract keywords Jullien et al.
+(2024); Jin et al.
+(2024); Nievas et al.
 (2024).
-Yuan et al.
-(2023) prompt LLMs to augment eligibility criteria and pass the BERT-based embeddings of these criteria and patient data through a fully connected classification layer to predict patient-criterion eligibility.
-The trial-level prediction is then composed by asserting that all inclusion criteria and none of the exclusion criteria match.
-On six trials for stoke patients, the model achieved 0.91 F1 at criterion-level vs.
-0.81 F1 at trial-level.
-Criterion-Level Prediction
-Hamer et al.
-(2023) use OpenAI’s text-davince-003 model to first filter criteria that applies to a patient, then justify whether applicable criteria are met and to finally predict criterion-level labels based on previous explanations.
-This effectively reduces the total number of eligibility criteria (down to 300 from originally 4,100) a physician has to assess before recommending trials to a patient.
-Wornow et al.
-(2025) prompt LLMs with trial-patient pairs to generate structured JSON strings of criterion-level predictions, explanations with evidence and a self-declared confidence.
-The content of the prompt ranges from the entire patient record and a trial to a single criterion and relevant record snippets.
-GPT4 achieves a micro-F1 of 0.93 vs.
-0.91 points by the prior state-of-the-art on the 2018 N2C2 cohort dataset.
-The RECTIFIER Unlu et al.
-(2024) employs a RAG-based approach to find top-k relevant patient data stored as vector embeddings.
-The criteria and relevant records are prompted to the LLM to generate either Yes or No.
-Evaluated on just a single trial and a pool of 3000 private patient records, with clinical notes spanning three years, RECTIFIER achieved an accuracy higher than trained non-licensed staff overall (92.7% vs.
-89.1%), especially for the Symptomatic Heart Failure criterion (97.9% vs.
-91.7%).
-Trial Ranking
-One line of work investigates the role of neural re-rankers for ranking relevant trials.
-Pradeep et al.
-(2022) generate multiple independent queries from a patient description to retrieve relevant trials.
-The trial and patient description is then passed to a fine-tuned T5 model to generate a relevance label.
-Zhuang et al.
-(2023) propose a multi-stage pipeline that involves a hybrid sparse-dense retriever, initialized with PubmedBERT-base and fine-tuned to retrieve the top 1000 trials.
-A cross-encoder, initialized with PubmedBERT-large, is trained to re-rank these results.
-Finally, the top-20 trials are re-ranked using GPT4.
-Rybinski et al.
-(2024) leverage LLMs for query reformulation and trial re-ranking to outperform traditional BM25 and BERT-based retrieval models.
-They show that BM25 is an effective initial retriever, surpassed by BERT-based neural re-rankers.
+This is followed by retrieval, either using embeddings similarity only Lahiri et al.
+(2023); Richmond and Deshpande (2023); Saeidi et al.
+(2023); Ferber et al.
+(2024); Saeidi (2025) or a multi-stage retrieval with neural re-rankers Zhuang et al.
+(2023); Rybinski and Karimi (2023); Rybinski et al.
+(2024); Jin et al.
+(2024); Datta et al.
+(2025).
+Certain approaches re-rank the top trials using GPT models Zhuang et al.
+(2023); Rybinski et al.
+(2024); Datta et al.
+(2025), some prompt LLMs for relevance labels Pradeep et al.
+(2022), while others prompt LLMs to generate trial- or criterion-level eligibility labels Rybinski et al.
+(2024); Peikos (2023); Jin et al.
+(2024); Nievas et al.
+(2024); Jullien et al.
+(2024).
+Criterion-level labels are aggregated using set-based reasoning mechanisms Jullien et al.
+(2024), variations of the matching Equation 6.1 Jin et al.
+(2024); Nievas et al.
+(2024); Saeidi (2025), and by prompting LLMs Jin et al.
+(2024).
+7 Evaluation
+Standard IR metrics, such as, precision, recall, F1, precision at rank k, normalized discounted cumulative gains (NDCG@k) and mean reciprocal rank (MRR), are the most popular evaluation metrics.
+The 2018 N2C2 cohort selection and TREC CT benchmarks have high performance metrics (Table 1), which are further improved by newer methods.
+Notably, researchers in the medical domain have stressed the lack of consistent benchmark data, primarily, due to the dependence on manual review which also restricts the size of the data Kim and Quintana (2022); Kantor and Morzy (2024b).
+Besides accuracy metrics, Rybinski et al.
+(2024) report query latency, with GPT models having 15 times higher latency (47.8s) when compared to smaller pre-trained models for a 12% increase in NDCG@10.
 Jin et al.
-(2024) use GPT4 to first extract query keywords to retrieve top-k trials and then prompt GPT4 with patient data and eligibility criteria to generate criterion-level labels and explanations.
-They use a combination of manually-defined and GPT4 prompted scores to determine a trial-level score for final raning.
-Their TrialGPT performs better on the TREC datasets compared to smaller domain-specific transformer models and show a screening time reduction by 42%.
-Trial-LLama by Nievas et al.
-(2024) extend this study by extensive fine-tuning of open-source LLaMA models.
-6 Discussion
-Annotated Corpora
-There have been few attempts to analyze and structure the clinical trials repository.
-As a result, finding similar trials and analyzing systematic failure cases is notoriously difficult.
-Chia, a manually annotated corpus of 1000 clinical trials created by Kury et al.
-(2020) has criterion-level annotations of entities, relations and Boolean logic graphs.
-Such annotations on trial-patient matching benchmarks can help us measure model fairness.
-Scaling patient data collections is also challenging as evident from Table 1, where patient records are in the hundreds.
-Individual records, curated manually, range from keywords via short patient summaries to documents over a time period.
-The format ranges from free text via semi-structured in-line tables to full-fledged databases.
-The rich and large MIMIC dataset, remains largely untapped for the task of trial-patient matching.
-LLMs provide a promising avenue for solving the automatic annotation at scale.
-Grouth-truth Labels
-Access to trial-patient eligibility labels, i.e., {eligible, ineligible, not-relevant}, is integral for training and evaluation.
-As they are obtained from human experts, they are difficult to scale and take time.
-Hence, many systems are evaluated retrospectively on patient data from private repositories such as clinics that maintain participation records in past trials Wong et al.
-(2023); Yuan et al.
-(2023); Hamer et al.
-(2023); Unlu et al.
-(2024).
-The largest public annotations from the TREC CT tracks have an average of 700 trial annotations per patient description.
-Bedi et al.
-(2025) found that of the 519 studies between 2022 and 2024 that evaluated LLM applications in healthcare, only 5% used real patient data.
-Jointly, the major public data sources, e.g., ClinicalTrials.gov and MIMIC datasets, present an exciting opportunity to develop generalizable matching methods.
-Formalization
-We formalize the task of trial-patient matching in Equation 3 through simple set operations.
-Surprisingly, except Yuan et al.
-(2023), none of the related work explicitly formalize the task.
-This coincides with a lack of aggregation strategies at the criterion-level Hamer et al.
-(2023); Wornow et al.
-(2025); Unlu et al.
-(2024).
-Formalizing the task, guides Yuan et al.
-(2023) to design a loss function that accounts for the contrastive requirements of inclusion and exclusion criteria.
-The criterion-level labels by Yuan et al.
-(2023), while not being directly interpretable, are aggregated as in Equation 3 for the trial-level labels.
-Wong et al.
-(2023) could potentially explain prediction via logical component mismatches, though the current F1 measures are below 50 points.
-Clinical experts, who analyze matches for the final cohort selection, strictly follow Equation 3.
-Criterion-level predictions and explanations have shown to reduce screening time by 42% Jin et al.
-(2024).
-Core Limitations of LLMs
-According to Harrer (2023) the core limitations affecting LLM adoption are unfiltered pre-training, which does not differentiate between facts, opinions, or misinformation; lack of self-assessment, where a model generates invalid but syntactically and semantically coherent sentences; non-determinism: where surface-form prompt variations lead to drastic changes in the output and repeatability is not guaranteed under consistent input conditions; knowledge recall: where updating outdated data or injecting new information requires expensive retraining since the mechanisms of memory in LLMs are not well understood.
-In Section 7, we discuss how knowledge recall and lack of self-assessment surface through reasoning errors.
-These limitations pose a direct challenge to the transparency and accountability principles of AI for health laid down by the World Health Organization Guidance (2021).
-Societal Impact
-Recent research measured distinct bias in disease diagnosis across gender, age and disease in popular LLM models (GPT4, ChatGPT and Qwen) Zhao et al.
-(2024b) and found that GPT4 tends to stereotype demographic presentations when generating diagnoses Zack et al.
-(2024).
-There is a lack of such analyses and mitigation methods in trial-patient matching and a lack of diverse datasets.
-The popularity of propreitary models for biomedical tasks Lu et al.
-(2024) including trial-patient matching Wornow et al.
-(2025); Jin et al.
-(2024) further ethical concerns on handling sensitive patient data.
-Alber et al.
-(2025) show that LLMs are prone to making medical misjudgments by replacing just 0.001% of the training data with medical misinformation.
-More research is required to determine the feasibility of locally deploying open-source models, the associated costs of maintaining them and keeping them up-to-date.
+(2024) report a 42% reduction in screening time when using LLM-generated predictions and explanations.
+The measurement of human preference of LLM generations, bias in model predictions and the informativeness of LLM explanations ranges from limited to non-existent.
 Figure 2: Taxonomy of errors in LLM-generations.
-7 Taxonomy of LLM Errors
-Since there is no one accepted taxonomy of errors, we come across semantically overlapping error types that are used to evaluate LLM generated explanations.
-For instance, incorrect reasoning and lack of knowledge are recognized as independent error types in Jin et al.
+7.1 Taxonomy of LLM Errors
+Since there is no one accepted taxonomy of errors, we often come across inconsistent and semantically overlapping categories of errors in LLM generated explanations.
+For e.g., incorrect reasoning and lack of knowledge are recognized as independent error types by Jin et al.
 (2024); Hamer et al.
 (2023); Nievas et al.
-(2024), even though the latter often leads to the former.
+(2024), even though the latter often leads to the former (full list in Appendix B).
 In the proposal by Liévin et al.
 (2024), reasoning errors are separate from reading comprehension, even though the instances of incorrect reading comprehension occur when the model ignores contextual information and “reasons” using its learned knowledge.
-We therefore identify where errors occur and break them down further by an identifiable source of the error (Figure 2).
+Here, we identify where errors occur and break them down further by an identifiable source of the error (see Figure 2).
 Information Isolation Errors
@@
 Reasoning Errors
-Any error in the explanations provided by LLMs fall into this category.
-The most common source is insufficient explicit data.
-Basically, LLMs fail to draw logical conclusions from given data, while an expert can.
+An error in the explanations provided by LLMs falls into this category.
+The most common source is insufficient explicit data, which occurs when LLMs fail to draw logical conclusions from given data, while an expert can.
 This stems from previously unseen data (“lack of knowledge”) or the inability to recall prior information (“lack of implicit reasoning”) or the inability to infer from context (“insufficient data”).
@@
 (2024)).
-The opposite situation, when a prediction is correct despite an incorrect explanation, is more difficult to report unless explanations for correct predictions are evaluated as well.
-It may seem as if the LLM is taking a shortcut, hinting towards a bias that the model picked up during training.
+The opposite situation, when a prediction is correct despite an incorrect explanation, is more difficult to evaluate.
+This shortcut, similar to cognitive biases in humans, hints towards a bias that the model picked up during training.
 Both these cases negatively impact transparency and accountability of the matching system.
+8 Discussion
+Annotated Corpora and their Size
+Despite attempts to structure clinical trials Chen et al.
+(2022), finding similar trials and analyzing systematic failure cases is notoriously difficult Rybinski et al.
+(2020); White et al.
+(2023).
+Criterion-level annotations require significant manual annotations, thereby limiting the size of the corpus.
+For instance, Chia Kury et al.
+(2020) and LCT Dobbins et al.
+(2022), are manually annotated corpora of 1000 trials each, while supervised annotators such as, EliIE Kang et al.
+(2017) and Criteria2Query Yuan et al.
+(2019), require expensive manual labels (230 disease-specific clinical trials in this case).
+Eligibility labels on real patient data from real enrollment status, necessary to discount reviewer bias, are only available on a small-scale, if at all, due to the trade-off between scale and privacy Kim and Quintana (2022).
+Systems thus evaluated on private retrospective data Wong et al.
+(2023); Yuan et al.
+(2023); Unlu et al.
+(2024) cannot be transparently compared.Kantor and Morzy (2024b) stress on standardized benchmarks as the dependence on manual evaluation hinders meta-analyses and comparison between different studies.
+Public annotations, such as the TREC CT tracks have an average of 700 trial annotations per patient for less than 100 patients, while the N2C2 has only a few thousands of criterion-level annotations.
+Jointly, the major public corpora, e.g., ClinicalTrials.gov and MIMIC datasets, present an opportunity to build on the proposal by Kim and Quintana (2022) to generate large-scale data using automated methods.
+Dimensions of Evaluation
+The benchmarks in Table 1 focus on model accuracy, corroborating the result from Bedi et al.
+(2025), which reports that more than 95% of studies use accuracy as the primary dimension of evaluation, while fairness, bias and uncertainty are less frequently measured.
+Omar et al.
+(2024) reviewed 27 clinical trials evaluating LLMs in healthcare also found the accuracy and reliability standards for LLM use to be undefined.
+Further results from Nemati et al.
+(2025), who benchmarked the annotation ability of LLMs across 9 performance metrics, show that while LLMs consistently score high on precision, recall and F1 (lowest being 0.8), their scores highly vary on semantic similarity, factual consistency, relevance, fluency, consistency and coherence (ranging from 0.1 to 0.9) highlighting the need for multiple dimensions of evaluation.
+Formalization
+Equation 6.1 highlights the importance of aggregation and priority.
+Surprisingly, very few works explicitly formalize the matching task.
+This lack of formalization coincides with an absence of aggregation strategies for trial-level predictions Hamer et al.
+(2023); Wornow et al.
+(2025); Unlu et al.
+(2024) and others.
+Formalization guides Yuan et al.
+(2023) to design a loss function that accounts for the contrastive requirements of inclusion and exclusion criteria, Jullien et al.
+(2024) and Saeidi (2025) to define re-ranking scores.
+Societal Impact
+Recent research measured distinct bias in disease diagnosis across gender, age and disease in popular LLM models (GPT4, ChatGPT and Qwen) Zhao et al.
+(2024b) and found that GPT4 tends to stereotype demographic presentations when generating diagnoses Zack et al.
+(2024).
+Alber et al.
+(2025) show that LLMs are prone to making medical misjudgments by replacing just 0.001% of the training data with medical misinformation.
+All the LLM-based systems, except for the trial ranking models, are evaluated on a few disease-specific trials (see Table LABEL:tab:systems_overview_llm in Appendix C), the largest being 146 trials, covering 10 cancer types, evaluated by Hamer et al.
+(2023), making the generalizability of LLMs across diseases unclear.
+Data Sensitivity
+While we found several methods that deployed GPT models on Azure AI to comply with privacy regulations Unlu et al.
+(2024); Wong et al.
+(2023); Wornow et al.
+(2025), processing patient data with LLMs still raises serious ethical challenges due to the lack of HIPAA compliance Edemekong et al.
+(2024).
+As already discussed, creating large-scale realistic patient records while protecting their privacy is particularly challenging.
+There is interest in generating synthetic data, called digital twins, with limited access to real patient data as a viable privacy protected alternative Das et al.
+(2023); Wang et al.
+(2024).
+Core Limitations of LLMs
+According to Harrer (2023) the core limitations affecting LLM adoption are unfiltered pre-training, which does not differentiate between facts, opinions, or misinformation; lack of self-assessment, where a model generates invalid but syntactically and semantically coherent sentences; non-determinism: where surface-form prompt variations lead to drastic changes in the output and repeatability is not guaranteed under consistent input conditions; and, knowledge recall: where updating outdated data or injecting new information requires expensive retraining since the mechanisms of memory in LLMs are not well understood.
+In Section 7.1, we discussed how knowledge recall and lack of self-assessment surface through reasoning errors.
+These limitations pose a direct challenge to the transparency and accountability principles of AI for health laid down by the World Health Organization Guidance (2021).
 9 Future Directions
-Clinical trial recruitment spans multiple interactive parts.
-As interest in LLMs orchestrating an end-to-end pipeline and incorporating human interactions is gaining more attention Gao et al.
-(2024); Qiu et al.
-(2024), we focus on four promising directions.
+As interest in LLMs orchestrating an end-to-end pipeline and incorporating human interactions is gaining more attention Gao et al. (2024); Qiu et al. (2024), we focus on four promising directions.
 Trial Search
 Past trials that share a target population are important for designing new trials, recruiting patients, systematic reviews and meta-analyses.
-This task goes beyond keyword search to match eligibility criteria of two trials.
 This problem has seen little activity since clustering using lexical features Hao et al.
 (2014).
-Analogous systems, such as SemEHR Wu et al.
-(2018), exist for interactive semantic search over patient EHRs.
-Sun et al.
-(2022) combine multiple similarity measures to account for categorical features (study type, phase, and so on) and lexical features (condition name, intervention name, outcome measures, eligibility criteria) to find similar trials.
+Newer search methods include constructing a clinical trial knowledge graph Chen et al.
+(2022), searching via patient EHRs Wu et al.
+(2018) and designing features for similarity matchingSun et al.
+(2022).
 Jullien et al.
-(2023) deploy LLMs to find trials that match short descriptions using textual entailment and ranked evidences.
+(2023) use textual entailment in LLMs to find trials that match short descriptions.
 However, these models falter on inferences that require numerical reasoning and could not surpass a BM25 baseline for ranking evidence.
 Interactive Trial Design
-Investigators design trials independent of patient data.
-This creates a semantic gap between eligibility criteria and patient data.
-LLM agents have the potential of bridging this gap by suggesting data models underlying patient data for structuring eligibility criteria early on.
+LLM agents have the potential of bridging the semantic gap between eligibility criteria and patient data by suggesting data models underlying patient data for structuring eligibility criteria early on.
 This collaborative idea is not new Luo et al.
-(2013), however, designing eligibility criteria is a big challenge and has been handled post-hoc by optionally considering patient data models Kang et al.
+(2013), yet, designing eligibility criteria is a big challenge and has been handled post-hoc by optionally considering patient data models Kang et al.
 (2017); Sun and Loparo (2019); Liu et al.
@@
 Collaborative Trial Planning
-Eligibility criteria have a range of difficulty levels.
-A preemptive task can interact with patient records to estimate the size of the patient pool on the most restrictive criteria, say the presence of a rare genetic disorder.
-As patient records progressively follow uniform data models, it is more efficient to separate and structure criteria that require tabular operations (e.g., via structured queries) from criteria that require reasoning (temporal, numerical, negation), and to separate criteria that require expert analysis from being wrongly used in semantic searches Mozannar and Sontag (2020).
+Liu et al.
+(2025c) propose an iterative feature discovery model using LLM agents for interpretable trial outcome prediction.
 Markey et al.
-(2025) showed promising results on content relevance and suitability with room for improvement in logical reasoning and provenance when using LLMs to generate trial protocols.
-Training LLM agents to identify and distribute tasks and then aggregating is a fertile ground for innovation.
+(2025) showed promising results on content relevance and suitability of trial protocols generated using LLMs, with room for improvement in logical reasoning and provenance.
+Similar to Shi et al.
+(2024), who propose collaboration of agents for knowledge-augmentation and reasoning, LLM agents can identify and distribute tasks and aggregate them to a final result.
+Another important operation is learning to defer to experts Mozannar and Sontag (2020), which can separate operable criteria, such as those that require tabular operations (e.g., via structured queries) or reasoning (temporal, numerical, negation), from criteria that require expert feedback.
 Explainable Matches
-In the context of LLMs, accurately explaining black-box predictions in human-understandable form becomes very challenging Zhao et al.
+Explaining black-box LLM predictions in human-understandable form is very challenging Zhao et al.
 (2024a).
-Explainability research in clinical trial matches has been limited to prompting LLMs and manually evaluating chain-of-thought generations post-hoc, which is only one of the many facets of evaluating explainability Nauta et al.
+Explainability in clinical trial matches is limited to chain-of-thought generations, which is only one of the many facets of explainability Nauta et al.
 (2023); Bodria et al.
 (2023); Chen and Eickhoff (2024).
-Moreover, it is unclear how helpful these explanations are, especially, as discussed in Section 7, since they give rise to inconsistency errors and are categorized into ambiguous error labels.
+Wong et al.
+(2023) could potentially explain eligibility prediction via logical component (mis)matches.
+Furthermore, the utility of these explanations are unclear, as they have been only evaluated manually.
+We hope that the error taxonomy discussed in Section 7.1 assists in systematic evaluation of explanations.
 10 Conclusion
-The ability of generative LLMs to generalize comes with a trade-off between reasoning and consistency.
-This work is the first to survey the rapidly evolving role of generative LLMs in clinical trial recruitment.
-We provide a modern perspective on the challenges in adopting LLMs to clinical research and point out promising directions for future work.
+The task of clinical trial recruitment, that matches patients to a clinical trial via its participation eligibility criteria, benefits from knowledge aggregation and reasoning abilities of LLMs.
+In this survey, we critically examine the evolving role of LLM technologies in clinical research.
+We analyze the main components in a clinical trial recruitment process and provide a modern perspective on the challenges in adopting LLMs to clinical research, such as the size of the benchmarks, the dimensions of evaluation and data sensitivity.
+We hope that this serves as a valuable resource for future research.
 11 Limitations
@@
 Given that this is a rapidly evolving field, we have made our best effort to include a comprehensive view of available resources and methods.
-It is possible that more sophisticated methods using the latest technology already exist (e.g., in the form of proprietary products), but are not yet made public.
-Appendix A Trial-Patient Matching Systems
-We show all the classical systems in Table 2 and the LLM-based systems in Table LABEL:tab:systems_overview_llm, by their task, method, data source characteristics and limitations.
-Appendix B Systematic Reviews versus Surveys
-A systematic review is a common practice in the medical research community with standardized reporting guidelines.
-The Preferred Reporting Items for Systematic Reviews and Meta-Analyses (PRISMA) Vrabel (2015) is the most widely used guideline.
-Just as finding patients given a trial eligibility criteria, this guideline requires strict inclusion and exclusion criteria to be applied to scientific literature to be included in a systematic review.
-In contrast, the guidelines for writing surveys in our community, for instance, as outlined for authors by ACM Computing Surveys , TACL and ARR aim to draw perspectives on an evolving topic of interest.
+It is possible that more sophisticated methods using the latest technology already exist (e.g., in the form of proprietary products), but are not yet made public or are only available as abstracts, as is common in some medical communities, for example, the Annual Meeting of the American Society of Clinical Oncology (ASCO).
+Appendix A Systematic Reviews versus Surveys
+A systematic review is a common practice in the medical research community, with standardized reporting guidelines Moher et al.
+(2009); Page et al.
+(2021).
+This guideline has a checklist of required items in the title, abstract, introduction, methods, results, discussion and other information that every study needs to fulfil.
+These studies are bound by strict inclusion and exclusion criteria applied to the scientific literature considered for any assessment.
+This rigor is necessary in evidence-based analysis of a specific research question.
+A survey is more flexible and provides a board coverage of the topic being discussed.
+For instance, the guidelines for writing surveys in our community as outlined by ACM Computing Surveys , TACL and ARR aim to draw perspectives on an evolving topic of interest.
 ACM Computing Surveys (long paper)
@@
 all papers are expected to include reviews of related literature. This category is meant for the papers that go beyond that, e.g. in scope or in establishing new interdisciplinary connections.
+Appendix B Reported Error Types
+In the following we illustrate the the error distributions by percentages as reported by previous studies.
+Errors reported in Jin et al. (2024) supplementary:
+  • Incorrect reasoning: 30.7%
+  • Lack of medical knowledge: 15.4%
+  • Ambiguous label definition: 26.9%
+  • other errors: 26.9% (self-conflicting)
+Errors reported in Hamer et al. (2023):
+  • Incorrect reading: 6.5%
+  • Insufficient knowledge: 2.2%
+  • Incorrect reasoning: 91.3%
+Errors reported in Nievas et al. (2024):
+  • Lack of knowledge  55%
+  • Implicit criteria  15%
+  • Wrong reasoning
+  • Accurate reasoning, wrong decision
+  • Lack of restraint when expert opinion is needed
+  • Negated criteria error
+Appendix C Trial-Patient Matching Systems
+Table 3 gives an overview of all the classical systems in covered in this survey by their direction of approach, method, source data characteristics and limitations.
+Table LABEL:tab:systems_overview_llm covers all the LLM-based systems discussed in this survey by their task, method, data source characteristics and limitations.
+We notice the shift in the availability of patient sources towards public data in the LLM-based approaches.
+We also notice a shift towards trial ranking due to publicly available data.
 Direction Method Patient Source Trial Source Limitations Trial-centric Penberthy et al. (2010) Discrete-data filter and sub-word matching Availability:
@@
 Task Method Patient Source Trial Source Limitations
+Criterion-level prediction Hamer et al. (2023)
+1-shot prompt to LLM for criteria level prediction with explanation
+Availability: Private (synthetic patient profile)
+Size: 10
+Mode: Short text
+Type: Specific
+Availability: Mixed (clinicalTrials.gov, EudraCT)
+Size: 146 clinical trials
+Type: Specific
+Additional Annotation: Expert review
+Majority of criterion-level errors due to incorrect reasoning (91%)
+Unlu et al. (2024)
+RAG-based
+Availability: Private
+Size: 2,276
+Mode: Tables; text
+Availability: Public∗∗
+Size: 1
+Type: Specific
+RAG pipeline not published. Evaluated on a single trial.
+Ferber et al. (2024)
+Criterion-level boolean eligibility labels
+Availability: Private
+Size: 51
+Mode: EHR
+Availability: Public∗∗
+Size: 105,600
+Type: Specific
+Use of proprietary model limits evaluation on synthetic patient data.
+Beattie et al. (2024)
+Vector database of patient notes. Prompt (criteria, top-k patient notes) pair to LLM
+Availability: Public (2018 N2C2: Cohort Selection) Some evaluation is on test subset (40 of 182).
+Expert guidance requires manual expertise for every criterion.
+Wornow et al. (2025)
+Vector database of patient notes. Prompt (criteria, top-k patient notes) pair to LLM
+Availability: Public (2018 N2C2: Cohort Selection) Doesn’t consider exclusion criteria
+67% incorrect decision despite correct reasoning.
+Saeidi (2025)
+Prompt LLM with a fine-tuned BERT-based concept embeddings from patient and criteria to predict binary eligibility labels.
+Availability: Public (2018 N2C2: Cohort Selection) LLM reasoning capabilities not explored.
 Trial-level prediction Wong et al. (2023)
@@
 High variance between different trials: F1 range: 0.39 - 0.96.
-Criterion-level prediction Hamer et al. (2023)
-1-shot prompt to LLM for criteria level prediction with explanation
-Availability: Private (synthetic patient profile)
-Size: 10
-Mode: Short text
-Type: Specific
-Availability: Mixed (clinicalTrials.gov, EudraCT)
-Size: 146 clinical trials
-Type: Specific
-Additional Annotation: Expert review
-Majority of criterion-level errors due to incorrect reasoning (91%)
-Unlu et al. (2024)
-RAG-based
-Availability: Private
-Size: 2,276
-Mode: Tables; text
-Availability: Public∗∗
-Size: 1
-Type: Specific
-RAG pipeline not published. Evaluated on a single trial.
-Wornow et al. (2025)
-Vector database of patient notes
-Prompt (criteria, top-k patient notes) pair to LLM
-Availability: Public (2018 N2C2: Cohort Selection) Evaluated on one trial
-Doesn’t consider exclusion criteria
-67% incorrect decision despite correct reasoning.
 Trial ranking Pradeep et al. (2022)
@@
 Validation on textual patient note does not transfer well to testing on structured patient note
-Rybinski et al. (2024)
+Cosine similarity between patient and trial embeddings obtained using GPT by Richmond and Deshpande (2023) and using Sentence Transformer by Lahiri et al.
+(2023).
+Availability: Public (TREC CT 2023) Very low MAP of 0.02 Richmond and Deshpande (2023) and close to zero for Lahiri et al.
+(2023)
+Kusa et al. (2023b)
+Sentence query formulation using GPT-3.5. Query enrichment using Kusa et al. (2023a). Trial-level prediction of re-ranked pairs using GPT-3.5.
+Availability: Public (TREC CT 2023) Zero-shot LLM prompts provide marginal improvement over neural-rerankers
+Peikos (2023)
+Utilize GPT3.5 to obtain trial-level labels for final re-ranking on top of lexical and neural re-rankers
+Availability: Public (TREC CT 2023) Query processing is template-based and excludes negative information.
+Ferber et al. (2024)
+No-SQL query formulation from patient EHR using GPT-4o for initial retrieval, followed by vector embedding re-ranking.
+Availability: Private
+Size: 51
+Mode: EHR
+Availability: Public∗∗
+Size: 105,600
+Type: Specific
+Use of proprietary model limits evaluation on synthetic patient data.
+Rybinski and Karimi (2023); Rybinski et al. (2024)
 Multi-stage retriever using LLMs
 Availability: Public (TREC CT 2021, 2022, 2023) High latency of GPT4 vs a small performance boost
+Jullien et al. (2024)
+LLM-guided basic attribute extraction for trial retrieval and filtering. Criterion-level LLM predictions fed to set-reasoning-based re-rank scoring functions.
+Availability: Public (TREC CT 2022) LLMs underperform in exclusion criteria labelling.
 Jin et al. (2024)
@@
 Criterion-level LLM prediction aggregated to trial scores
-Availability: Public (Koopman and Zuccon (2016), TREC CT 2021 & 2022)
-Size: 58; 75; 50
-Mode: Short text
-Additional annotations: 3,621; 26,149; 26,581 trial-patient matches
-Availability: Public∗∗
-Size: 375,580
+Availability: Public (Koopman and Zuccon (2016), TREC CT 2021, 2022)
 Additional annotation: 1,015 criterion-patient labels
@@
 Significantly high fine-tuning costs
+Datta et al. (2025)
+Lexical retrievers use LLM generated query expansion. LLM predicts trial-level label and a criterion-level rationale with a matching score. Final ranking based on matching score.
+Availability: Public (TREC CT 2023) Trials prefiltered by disorder-specific keywords.
+Manually curated disorder-specific instructions.
+Model predicted trial-level label is not evaluated
+Saeidi et al.
+(2023); Saeidi (2025) Embed patients and trials to concept vector space.
+Use variations of Equation 6.1 to compute relevance scores.
+Availability: Public (TREC CT 2023) No direct connection between criterion-level LLM predictions and trial-level relevance score computations
 EudraCT (European Union Drug Regulating Authorities Clinical Trials) is the European clinical trials database EudraCT .

### v3

@@
 Selecting the precise and required size of patient population is crucial for trial completion.
-According to various estimates, more than 50% of aborted clinical trials fail due to low accrual rates, and 80% of all clinical trials do not manage to recruit the required patient cohorts within the allotted time Clinical Trials Arena (2012); Williams et al.
-(2015); Pharmaceutical Technology (2019).
-While the trend has steadily declined over the past decade, with the intensive use of technology-aided solutions, efficient patient recruitment remains the most crucial bottleneck in clinical trial research Clinical Trials Arena (2022).
-As electronic health records (EHRs) of patients become more accessible, clinical researchers adopt machine intelligence and develop explainable systems to correctly interpret model predictions Murdoch and Detsky (2013); Payrovnaziri et al.
-(2020); von Itzstein et al.
-(2021).
-There are methods leveraging LLMs in clinical trial research for cohort retrieval and modelling Fang et al.
-(2022); Tian et al.
-(2023); Park et al.
-(2024); Liu et al.
-(2025a); Wang et al.
-(2025), trial design Reinisch et al.
-(2024); Curran et al.
-(2024); Bornet et al.
-(2025); Neehal et al.
-(2025), trial search White et al.
-(2023); Rybinski et al.
-(2020), trial matching Jin et al.
-(2024); Nievas et al.
-(2024); Wornow et al.
-(2025), predicting trial outcomes and duration Reinisch et al.
-(2024); Yue et al.
-(2024a, b); Liu et al.
-(2025c), assessing risk of bias Lai et al.
-(2024); Ji et al.
-(2025), and extracting clinical trial results Lee et al.
-(2024), even as the community is coming up with recommended practices for responsible use of AI in the entire drug development process Geraci et al.
-(2025).
+According to various estimates, more than 50% of aborted clinical trials fail due to low accrual rates, and 80% of all clinical trials do not manage to recruit the required patient cohorts within the allotted time clinicaltrialarena2012; williams2015terminated; pharmaceutical2019.
+While the trend has steadily declined over the past decade, with the intensive use of technology-aided solutions, efficient patient recruitment remains the most crucial bottleneck in clinical trial research clinicaltrialarena2022.
+As electronic health records (EHRs) of patients become more accessible, clinical researchers adopt machine intelligence and develop explainable systems to correctly interpret model predictions murdoch2013inevitable; payrovnaziri2020explainable; vonitzstein2021application.
+There are methods leveraging LLMs in clinical trial research for cohort retrieval and modelling fang2022combining; tian2023parsing; park2024criteria2query; liu2025neuralcohort; wang2025trialpanorama, trial design reinisch2024ctp; curran2024examining; bornet2025analysis; neehal2025large, trial search white2023clinidigest; rybinski2020clinical, trial matching jin2024matching; nievas2024distilling; wornow2025zero, predicting trial outcomes and duration reinisch2024ctp; yue2024clinicalagent; yue2024trialdura; liu2025autoct, assessing risk of bias lai2024assessing; ji2025robguard, and extracting clinical trial results lee2024seetrials, even as the community is coming up with recommended practices for responsible use of AI in the entire drug development process geraci2025current.
 Figure 1: Components in a patient recruitment process: conventional linear flow (in orange) vs. our proposed LLM-assisted interactive flow (in purple).
 Figure 1 shows the components in clinical trial recruitment, namely, data sourcing, information extraction, matching, and evaluation.
-An expert reviews several hundred patients per trial and can end up spending hours on one patient, hence incurring significant costs Penberthy et al.
-(2012); Ni et al.
-(2015).
-Even simple automation using table queries and lexical searches saves between 165 hours to 1,329 hours of reviewing time when compared to manual evaluation Penberthy et al.
-(2010).
-In the past, the patient recruitment process has seen relatively low adoption of the pretrained language models He et al.
-(2020); Harrer (2023); Lu et al.
-(2024).
-Generative LLMs serve as knowledge aggregators, and through their reasoning and instruction-following capabilities, they have revived research in the task of trial and patient matching Jin et al.
-(2024); Nievas et al.
-(2024); Rybinski et al.
-(2024); Wornow et al.
-(2025).
+An expert reviews several hundred patients per trial and can end up spending hours on one patient, hence incurring significant costs penberthy2012effort; ni2015increasing.
+Even simple automation using table queries and lexical searches saves between 165 hours to 1,329 hours of reviewing time when compared to manual evaluation penberthy2010automated.
+In the past, the patient recruitment process has seen relatively low adoption of the pretrained language models he2020clinical; harrer2023attention; lu2024large.
+Generative LLMs serve as knowledge aggregators, and through their reasoning and instruction-following capabilities, they have revived research in the task of trial and patient matching jin2024matching; nievas2024distilling; rybinski2024learning; wornow2025zero.
 Difference to Prior Work
 Despite the rapidly evolving landscape of LLM technology, there is no prior work surveying this area.
-Gueguen et al.
-(2025) evaluate public trial matching tools and Layne et al.
-(2025) compare the efficacy of open and proprietary LLM-assisted trial-patient matching in oncology.
-Systematic reviews in this topic, bound by strict selection criteria and highly specific research question, do not to capture the broad perspectives Kim and Quintana (2022); Idnay et al.
-(2022); Kantor and Morzy (2024b); Chen et al.
-(2025) (Appendix A differentiates between systematic reviews and surveys).
-Although it provides an overview of LLM approaches used in clinical trial matching, Chen et al.
-(2025) only briefly discusses associated challenges and future work.
+gueguen2025prospective evaluate public trial matching tools and layne2025large compare the efficacy of open and proprietary LLM-assisted trial-patient matching in oncology.
+Systematic reviews in this topic, bound by strict selection criteria and highly specific research question, do not to capture the broad perspectives kim2022review; idnay2022systematic; kantor2024machine; chen2025enhancing (Appendix A differentiates between systematic reviews and surveys).
+Although it provides an overview of LLM approaches used in clinical trial matching, chen2025enhancing only briefly discusses associated challenges and future work.
 Our Contribution
@@
 Also known as patient recruitment/enrollment/(pre-)screening, it is the process of matching patients (or a cohort) to a clinical trial via its eligibility criteria.
-Clinical trials have a dual nature, consisting of both universal and trial-specific requirements, making it challenging to design generalized approaches Idnay et al.
-(2023).
+Clinical trials have a dual nature, consisting of both universal and trial-specific requirements, making it challenging to design generalized approaches idnay2023clinical.
 This has traditionally resulted in linear trial-centric matching processes (illustrated in orange in Figure 1) with limited scope for interaction and feedback.
-Such systems generally have an initial data filter on the structured EHRs followed by keyword matches and concept identification Penberthy et al.
-(2010); Tun et al.
-(2023) or cohort-specific classifiers Zhang and Demner-Fushman (2017).
-Ni et al.
-(2015) represented both trials and patients as feature vectors, thus supporting retrievals for both trial and patient.
+Such systems generally have an initial data filter on the structured EHRs followed by keyword matches and concept identification penberthy2010automated; tun2023automatic or cohort-specific classifiers zhang2017automated.
+ni2015increasing represented both trials and patients as feature vectors, thus supporting retrievals for both trial and patient.
 Biomedical NLP and LLMs
-The biomedical NLP landscape is shifting from specialized pretrained language models, such as, BioBERT Lee et al.
-(2020), BioLM Lewis et al.
-(2020), PubmedBERT Gu et al.
-(2021), BioGPT Luo et al.
-(2022), MedCPT Jin et al.
-(2023), among many others Wang et al.
-(2023), towards instruction-following and chat-enabled LLMs (commonly termed as generative AI), used as is Nori et al.
-(2023); Kung et al.
-(2023) or fine-tuned for domain alignment, such as, Med-PaLM Singhal et al.
-(2023), Med-Alpaca Han et al.
-(2023) and LLaVA-Med Li et al.
-(2023).
-We point our reader to Thirunavukarasu et al.
-(2023); Liu et al.
-(2025b) for further reading.
-The Journal of American Medical Informatics Association (JAMIA), which published 41 articles on biomedical health and LLMs in a focus issue, also observed this shift towards generative AI Lu et al.
-(2024).
-Among these, the OpenAI Achiam et al.
-(2023) family of proprietary models (GPT3.5 and GPT4 being the most common) is used far more often than open-sourced models Touvron et al.
-(2023); Bai et al.
-(2023); Jiang et al.
-(2023) models.
-Further, fine-tuning generative LLMs on the biomedical domain offers limited performance gain Dorfner et al.
-(2024).
-In fact, Alber et al.
-(2025) show that models are more prone to propagate medical misinformation encountered during fine-tuning despite performing well on benchmarks.
-Finally, Bedi et al.
-(2025) report that real patient data is used in less than 5% of the 519 biomedical studies using LLMs, and fairness, bias, uncertainty and deployment considerations are rarely assessed.
+The biomedical NLP landscape is shifting from specialized pretrained language models, such as, BioBERT lee2020biobert, BioLM lewis2020pretrained, PubmedBERT gu2021domain, BioGPT luo2022biogpt, MedCPT jin2023medcpt, among many others wang2023pre, towards instruction-following and chat-enabled LLMs (commonly termed as generative AI), used as is nori2023capabilities; kung2023performance or fine-tuned for domain alignment, such as, Med-PaLM singhal2023large, Med-Alpaca han2023medalpaca and LLaVA-Med li2023llava.
+We point our reader to thirunavukarasu2023large; liu2025application for further reading.
+The Journal of American Medical Informatics Association (JAMIA), which published 41 articles on biomedical health and LLMs in a focus issue, also observed this shift towards generative AI lu2024large.
+Among these, the OpenAI achiam2023gpt family of proprietary models (GPT3.5 and GPT4 being the most common) is used far more often than open-sourced models touvron2023llama; bai2023qwen; jiang2023mistral models.
+Further, fine-tuning generative LLMs on the biomedical domain offers limited performance gain dorfner2024biomedical.
+In fact, alber2025medical show that models are more prone to propagate medical misinformation encountered during fine-tuning despite performing well on benchmarks.
+Finally, bedi2025testing report that real patient data is used in less than 5% of the 519 biomedical studies using LLMs, and fairness, bias, uncertainty and deployment considerations are rarely assessed.
 State of Adoption of NLP Advancements
-Kantor and Morzy (2024b)’s study on AI adoption rates in eligibility criteria parsing, report very low adoption rates of generative AI, with BERT family of models being most popular and generative models coming into the picture only since 2024.
-A systematic review of the role of NLP systems in patient recruitment identified only 11 studies Idnay et al.
-(2022).
-Heterogeneous outcomes, diverse results, a dependence on small retrospective data and a lack of common standardized benchmarks drive the gap in NLP research and their adoption in real-world settings Idnay et al.
-(2024); Kantor and Morzy (2024b).
-A study by Idnay et al.
-(2023), investigating how clinical researchers screened patients, highlights the challenges of universal and domain-specific nature of the eligibility criteria and makes recommendations to build interactive, flexible and transparent recruitment strategies.
-Interestingly, when Corbaux et al.
-(2024) categorize tools for oncological trial matching, they indicate that the automatic methods still fall in the research and development phase, yet to be commercially available.
+kantor2024machine’s study on AI adoption rates in eligibility criteria parsing, report very low adoption rates of generative AI, with BERT family of models being most popular and generative models coming into the picture only since 2024.
+A systematic review of the role of NLP systems in patient recruitment identified only 11 studies idnay2022systematic.
+Heterogeneous outcomes, diverse results, a dependence on small retrospective data and a lack of common standardized benchmarks drive the gap in NLP research and their adoption in real-world settings idnay2024clinical; kantor2024machine.
+A study by idnay2023clinical, investigating how clinical researchers screened patients, highlights the challenges of universal and domain-specific nature of the eligibility criteria and makes recommendations to build interactive, flexible and transparent recruitment strategies.
+Interestingly, when corbaux2024patients categorize tools for oncological trial matching, they indicate that the automatic methods still fall in the research and development phase, yet to be commercially available.
 3 Methodology
 We queried Google Scholar with the keywords “clinical trial”, “cohort discovery”, “patient recruitment”, “trial recruitment”, “trial matching” in conjunction with “llm”, “language model”, “gpt” and inspected the first ten result pages, between the years 2019 and 2025 (both inclusive).
-We also used Jin et al.
-(2024) and Wornow et al.
-(2025) as seeds and recursively traced their citations to efficiently captured the evolution of the predictive methods in this field.
+We also used jin2024matching and wornow2025zero as seeds and recursively traced their citations to efficiently captured the evolution of the predictive methods in this field.
 We organized papers that described LLM-based approaches and corpora for clinical trial recruitment into data sourcing (Section 2), information extraction and parsing (Section 5), trial-patient matching (Section 6) and evaluation (Section 7).
@@
 2018 N2C2
-Stubbs et al. (2019)
+stubbs2019cohort
 #patients = 288
@@
 ranking
-Koopman and Zuccon (2016) #patient summaries (22 words avg) = 60
+koopman2016test #patient summaries (22 words avg) = 60
 #patient description (78 words avg) = 60
@@
 TREC CT 2021
-Soboroff (2021)
+soboroff2021overview
 #patient descriptions = 75
@@
 TREC CT 2022
-Roberts et al. (2022)
+roberts2022overview
 #patient descriptions = 50
@@
 TREC CT 2023
-Rybinski et al. (2024)∗∗
+rybinski2024learning∗∗
 #patient tables = 40
@@
 No participants, since this was not a challenge.
-∗∗Borrowed from Rybinski et al.
-(2024) as TREC CT 2023 does not have a published track overview.
+∗∗Borrowed from rybinski2024learning as TREC CT 2023 does not have a published track overview.
 Table 1: Overview of the public trial-patient matching benchmarks.
@@
 We analyze five trial-patient matching benchmarks.
-  • 2018 N2C2 Cohort Selection Stubbs et al. (2019) for criterion-level eligibility prediction. [^1] Currently unavailable as of 2024 Nov 6.
-  • Koopman and Zuccon (2016) for ranking trials.
+  • 2018 N2C2 Cohort Selection stubbs2019cohort for criterion-level eligibility prediction. [^1] Currently unavailable as of 2024 Nov 6.
+  • koopman2016test for ranking trials.
   • Text REtrieval Conference Clinical Trial (TREC CT) [^2] Available at https://www.trec-cds.org/. tracks 2021, 2022, 2023 for ranking trials.
@@
 4.1 Trial Data
-The TREC CT benchmarks and Koopman and Zuccon (2016) source the ClinicalTrials.gov registry.
+The TREC CT benchmarks and koopman2016test source the ClinicalTrials.gov registry.
 Meanwhile, the N2C2 benchmark focuses on a single trial.
@@
 The N2C2 benchmark contains 288 de-identified longitudinal records of patients.
-The trial ranking benchmarks use up to 75 synthetic patient profiles, comprising keywords defining cohorts Koopman and Zuccon (2016), short textual patient descriptions (TREC CT 2021, 2022) and questionnaire templates (TREC CT 2023).
-The MIMIC database Johnson et al.
-(2016, 2020) is the largest anonymized public database of structured and unstructured data of 299,712 patients.
+The trial ranking benchmarks use up to 75 synthetic patient profiles, comprising keywords defining cohorts koopman2016test, short textual patient descriptions (TREC CT 2021, 2022) and questionnaire templates (TREC CT 2023).
+The MIMIC database johnson2016mimic; johnson2020mimic is the largest anonymized public database of structured and unstructured data of 299,712 patients.
 While this dataset is popular for training biomedical LLMs (as mentioned in Section 2), it is not used in any of the five benchmarks, possibly due to significant challenges in obtaining eligibility labels at this scale.
@@
 The TREC CT annotations were created by pooling the top-k results from all participating teams.
-Medical experts manually annotated this pool of trial-patient matches Koopman and Zuccon (2016); Soboroff (2021); Roberts et al.
-(2022).
-Out of a total of 35,394 pooled trial-patient matches in the 2022 edition, 11% were judged as eligible, 9% as ineligible and 80% as not relevant with an average of 700 trials judged per patient Roberts et al.
-(2022).
-The relevance judgments were more balanced between the three labels Rybinski et al.
-(2024) in TREC CT 2023.
-All benchmarks depend on manual annotation from experts, which is time-consuming and challenging to scale Kim and Quintana (2022).
+Medical experts manually annotated this pool of trial-patient matches koopman2016test; soboroff2021overview; roberts2022overview.
+Out of a total of 35,394 pooled trial-patient matches in the 2022 edition, 11% were judged as eligible, 9% as ineligible and 80% as not relevant with an average of 700 trials judged per patient roberts2022overview.
+The relevance judgments were more balanced between the three labels rybinski2024learning in TREC CT 2023.
+All benchmarks depend on manual annotation from experts, which is time-consuming and challenging to scale kim2022review.
 5 Information Extraction and Parsing
-Extraction of medical entities evolved from a combination of rule-based heuristics and feature-based supervised sequence labelling models Kang et al.
-(2017); Yuan et al.
-(2019) via embedding-based neural models Khan et al.
-(2019); Tseo et al.
-(2020) to transformer-based pretrained models and LLMs Liu et al.
-(2021c); Zeng et al.
-(2020); Tian et al.
-(2021); Li et al.
-(2022); Murcia et al.
-(2024); Kantor and Morzy (2024a).
-Datta et al.
-(2024) use disease-specific prompting to extract structured information about entities and its attributes from criteria text.
-Gao et al.
-(2020); Zhang et al.
-(2020); Theodorou et al.
-(2023) use BERT-based embeddings to encode eligibility criteria and patient data.
-Patient data converted to search queries, via reformulation and expansion using LLMs, are particularly effective in trial retrieval Peikos et al.
-(2023); Rybinski et al.
-(2024); Peikos et al.
-(2024); Jin et al.
-(2024).
-Yuan et al.
-(2019); Tian et al.
-(2023); Park et al.
-(2024); Mugambi et al.
-(2024); Ziletti and D’Ambrosi (2025) use semantic parsing to translate eligibility criteria into logical forms ready for querying patient databases.
+Extraction of medical entities evolved from a combination of rule-based heuristics and feature-based supervised sequence labelling models kang2017eliie; yuan2019criteria2query via embedding-based neural models khan2019improving; tseo2020information to transformer-based pretrained models and LLMs liu2021clinical; zeng2020ensemble; tian2021transformer; li2022comparative; murcia2024automating; kantor2024finetuned.
+datta2024autocriteria use disease-specific prompting to extract structured information about entities and its attributes from criteria text.
+gao2020compose; zhang2020deepenroll; theodorou2023treement use BERT-based embeddings to encode eligibility criteria and patient data.
+Patient data converted to search queries, via reformulation and expansion using LLMs, are particularly effective in trial retrieval peikos2023investigating; rybinski2024learning; peikos2024leveraging; jin2024matching.
+yuan2019criteria2query; tian2023parsing; park2024criteria2query; mugambi2024leveraging; ziletti2025generating use semantic parsing to translate eligibility criteria into logical forms ready for querying patient databases.
 6 Trial-Patient Matching
@@
 Alternatively, a patient-centric approach, taken by a patient or their healthcare provider, matches relevant trials from a trial registry to a short description of the patient.
-Koopman and Zuccon (2016) and the TREC CT 2021, 2022, and 2023 are patient-centric benchmarks for ranking trials.
+koopman2016test and the TREC CT 2021, 2022, and 2023 are patient-centric benchmarks for ranking trials.
 Classical LLM-based !
@@
 6.2 LLM-based Approaches
-Unlike classical approaches Penberthy et al.
-(2010); Zhang and Demner-Fushman (2017); Yuan et al.
-(2019); Tun et al.
-(2023), LLM-based approaches do not rely on trial-specific heuristics.
+Unlike classical approaches penberthy2010automated; zhang2017automated; yuan2019criteria2query; tun2023automatic, LLM-based approaches do not rely on trial-specific heuristics.
 Table 2 lists the primary differences between the classical and LLM-based approaches (see Appendix C for a full comparison).
@@
 While criterion- and trial-level predictions are direction-agnostic, trial ranking is a patient-centric task.
-With enough patients, we could evaluate patient ranking, similar to the bidirectional implementation in Ni et al.
-(2015)’s work, though this has not yet been addressed.
+With enough patients, we could evaluate patient ranking, similar to the bidirectional implementation in ni2015increasing’s work, though this has not yet been addressed.
 Criterion-Level Prediction
 Here, methods utilize the reasoning capability of LLMs to obtain decision rationale and other context data in addition to the eligibility label.
-Hamer et al.
-(2023) use 1-shot prompt, where given the patient profile and the eligibility criteria, the LLM first labels each criterion as being applicable to the patient or not, followed by a list of rationales and finally, the eligibility labels for the applicable criteria.
-Unlu et al.
-(2024); Beattie et al.
-(2024) and Wornow et al.
-(2025) chunk longitudinal patient data and store them as vector embeddings.
-They prompt LLMs with criteria and relevant patient data chunks under zero-shot setting, with Beattie et al.
-(2024) providing expert criterion-level strategy in the prompts.
-Unlu et al.
-(2024) generate only the decision label from GPT4 and Beattie et al.
-(2024) and Wornow et al.
-(2025) generate criterion-level JSON objects, comprising the criteria label, the eligibility label and a rationale among other context.
-Both work with proprietary OpenAI models (GPT-3.5 and GPT-4) and Wornow et al.
-(2025) additionally used open-sourced Llama Jiang et al.
-(2023) and Mixtral models Jiang et al.
-(2024).
-Ferber et al.
-(2024) prompt GPT-4o to generate criterion-level boolean predictions to be reviewed by experts.
+hamer2023improving use 1-shot prompt, where given the patient profile and the eligibility criteria, the LLM first labels each criterion as being applicable to the patient or not, followed by a list of rationales and finally, the eligibility labels for the applicable criteria.
+unlu2024retrieval; beattie2024utilizing and wornow2025zero chunk longitudinal patient data and store them as vector embeddings.
+They prompt LLMs with criteria and relevant patient data chunks under zero-shot setting, with beattie2024utilizing providing expert criterion-level strategy in the prompts.
+unlu2024retrieval generate only the decision label from GPT4 and beattie2024utilizing and wornow2025zero generate criterion-level JSON objects, comprising the criteria label, the eligibility label and a rationale among other context.
+Both work with proprietary OpenAI models (GPT-3.5 and GPT-4) and wornow2025zero additionally used open-sourced Llama jiang2023mistral and Mixtral models jiang2024mixtral.
+ferber2024end prompt GPT-4o to generate criterion-level boolean predictions to be reviewed by experts.
 Trial-Level Prediction
-Wong et al.
-(2023) used GPT3.5 and GPT4 to extract and structure eligibility criteria into logical expression to be matched locally with structured patient information.
-Yuan et al.
-(2023) use LLMs to augment eligibility criteria and pass the BERT-based embeddings of these criteria and patient data through a fully connected classification layer to predict patient-criterion eligibility.
+wong2023scaling used GPT3.5 and GPT4 to extract and structure eligibility criteria into logical expression to be matched locally with structured patient information.
+yuan2023large use LLMs to augment eligibility criteria and pass the BERT-based embeddings of these criteria and patient data through a fully connected classification layer to predict patient-criterion eligibility.
 The trial-level prediction is then computed by asserting Equation 6.1.
@@
 Here, methods formulate effective queries, retrieve trials and re-rank them.
-Query processing techniques involve generating sentence queries from patient descriptions using a fine-tuned T5 model Pradeep et al.
-(2022) or zero-shot LLM prompts Saeidi et al.
-(2023); Kusa et al.
-(2023b), generating patient descriptions from trial data via 1-shot LLM prompts Zhuang et al.
-(2023), generating no-SQL queries via LLMs Ferber et al.
-(2024), using LLMs to reformulate and expand queries Rybinski et al.
-(2024); Peikos et al.
-(2024); Datta et al.
-(2025), and using LLMs to extract keywords Jullien et al.
-(2024); Jin et al.
-(2024); Nievas et al.
-(2024).
-This is followed by retrieval, either using embeddings similarity only Lahiri et al.
-(2023); Richmond and Deshpande (2023); Saeidi et al.
-(2023); Ferber et al.
-(2024); Saeidi (2025) or a multi-stage retrieval with neural re-rankers Zhuang et al.
-(2023); Rybinski and Karimi (2023); Rybinski et al.
-(2024); Jin et al.
-(2024); Datta et al.
-(2025).
-Certain approaches re-rank the top trials using GPT models Zhuang et al.
-(2023); Rybinski et al.
-(2024); Datta et al.
-(2025), some prompt LLMs for relevance labels Pradeep et al.
-(2022), while others prompt LLMs to generate trial- or criterion-level eligibility labels Rybinski et al.
-(2024); Peikos (2023); Jin et al.
-(2024); Nievas et al.
-(2024); Jullien et al.
-(2024).
-Criterion-level labels are aggregated using set-based reasoning mechanisms Jullien et al.
-(2024), variations of the matching Equation 6.1 Jin et al.
-(2024); Nievas et al.
-(2024); Saeidi (2025), and by prompting LLMs Jin et al.
-(2024).
+Query processing techniques involve generating sentence queries from patient descriptions using a fine-tuned T5 model pradeep2022neural or zero-shot LLM prompts saeidi2023malnis; kusa2023dossier, generating patient descriptions from trial data via 1-shot LLM prompts zhuang2023team, generating no-SQL queries via LLMs ferber2024end, using LLMs to reformulate and expand queries rybinski2024learning; peikos2024leveraging; datta2025patient2trial, and using LLMs to extract keywords jullien2024controlled; jin2024matching; nievas2024distilling.
+This is followed by retrieval, either using embeddings similarity only lahiri2023tmu; richmond2023leveraging; saeidi2023malnis; ferber2024end; saeidi2025streamlining or a multi-stage retrieval with neural re-rankers zhuang2023team; rybinski2023matching; rybinski2024learning; jin2024matching; datta2025patient2trial.
+Certain approaches re-rank the top trials using GPT models zhuang2023team; rybinski2024learning; datta2025patient2trial, some prompt LLMs for relevance labels pradeep2022neural, while others prompt LLMs to generate trial- or criterion-level eligibility labels rybinski2024learning; piekos2023unimib; jin2024matching; nievas2024distilling; jullien2024controlled.
+Criterion-level labels are aggregated using set-based reasoning mechanisms jullien2024controlled, variations of the matching Equation 6.1 jin2024matching; nievas2024distilling; saeidi2025streamlining, and by prompting LLMs jin2024matching.
 7 Evaluation
@@
 The 2018 N2C2 cohort selection and TREC CT benchmarks have high performance metrics (Table 1), which are further improved by newer methods.
-Notably, researchers in the medical domain have stressed the lack of consistent benchmark data, primarily, due to the dependence on manual review which also restricts the size of the data Kim and Quintana (2022); Kantor and Morzy (2024b).
-Besides accuracy metrics, Rybinski et al.
-(2024) report query latency, with GPT models having 15 times higher latency (47.8s) when compared to smaller pre-trained models for a 12% increase in NDCG@10.
-Jin et al.
-(2024) report a 42% reduction in screening time when using LLM-generated predictions and explanations.
+Notably, researchers in the medical domain have stressed the lack of consistent benchmark data, primarily, due to the dependence on manual review which also restricts the size of the data kim2022review; kantor2024machine.
+Besides accuracy metrics, rybinski2024learning report query latency, with GPT models having 15 times higher latency (47.8s) when compared to smaller pre-trained models for a 12% increase in NDCG@10.
+jin2024matching report a 42% reduction in screening time when using LLM-generated predictions and explanations.
 The measurement of human preference of LLM generations, bias in model predictions and the informativeness of LLM explanations ranges from limited to non-existent.
@@
 Since there is no one accepted taxonomy of errors, we often come across inconsistent and semantically overlapping categories of errors in LLM generated explanations.
-For e.g., incorrect reasoning and lack of knowledge are recognized as independent error types by Jin et al.
-(2024); Hamer et al.
-(2023); Nievas et al.
-(2024), even though the latter often leads to the former (full list in Appendix B).
-In the proposal by Liévin et al.
-(2024), reasoning errors are separate from reading comprehension, even though the instances of incorrect reading comprehension occur when the model ignores contextual information and “reasons” using its learned knowledge.
+For e.g., incorrect reasoning and lack of knowledge are recognized as independent error types by jin2024matching; hamer2023improving; nievas2024distilling, even though the latter often leads to the former (full list in Appendix B).
+In the proposal by lievin2024can, reasoning errors are separate from reading comprehension, even though the instances of incorrect reading comprehension occur when the model ignores contextual information and “reasons” using its learned knowledge.
 Here, we identify where errors occur and break them down further by an identifiable source of the error (see Figure 2).
@@
 This includes missed or incorrect NER, measurements (numeric or unit), temporal scopes and negations.
-LLMs are good at recognizing entities and measurements, but still struggle with negations Nievas et al.
-(2024).
+LLMs are good at recognizing entities and measurements, but still struggle with negations nievas2024distilling.
 Reasoning Errors
@@
 Inconsistency Errors
-Generating explanations to arrive at final answers unlock LLM’s reasoning capacities Wei et al.
-(2022).
+Generating explanations to arrive at final answers unlock LLM’s reasoning capacities wei2022chain.
 Even so, the explanation and the prediction can be inconsistent.
-LLMs may predict incorrectly despite correct explanation (reported as explanation-output mismatch in Nievas et al.
-(2024)).
+LLMs may predict incorrectly despite correct explanation (reported as explanation-output mismatch in nievas2024distilling).
 The opposite situation, when a prediction is correct despite an incorrect explanation, is more difficult to evaluate.
@@
 Annotated Corpora and their Size
-Despite attempts to structure clinical trials Chen et al.
-(2022), finding similar trials and analyzing systematic failure cases is notoriously difficult Rybinski et al.
-(2020); White et al.
-(2023).
+Despite attempts to structure clinical trials chen2022knowledge, finding similar trials and analyzing systematic failure cases is notoriously difficult rybinski2020clinical; white2023clinidigest.
 Criterion-level annotations require significant manual annotations, thereby limiting the size of the corpus.
-For instance, Chia Kury et al.
-(2020) and LCT Dobbins et al.
-(2022), are manually annotated corpora of 1000 trials each, while supervised annotators such as, EliIE Kang et al.
-(2017) and Criteria2Query Yuan et al.
-(2019), require expensive manual labels (230 disease-specific clinical trials in this case).
-Eligibility labels on real patient data from real enrollment status, necessary to discount reviewer bias, are only available on a small-scale, if at all, due to the trade-off between scale and privacy Kim and Quintana (2022).
-Systems thus evaluated on private retrospective data Wong et al.
-(2023); Yuan et al.
-(2023); Unlu et al.
-(2024) cannot be transparently compared.Kantor and Morzy (2024b) stress on standardized benchmarks as the dependence on manual evaluation hinders meta-analyses and comparison between different studies.
+For instance, Chia kury2020chia and LCT dobbins2022leaf, are manually annotated corpora of 1000 trials each, while supervised annotators such as, EliIE kang2017eliie and Criteria2Query yuan2019criteria2query, require expensive manual labels (230 disease-specific clinical trials in this case).
+Eligibility labels on real patient data from real enrollment status, necessary to discount reviewer bias, are only available on a small-scale, if at all, due to the trade-off between scale and privacy kim2022review.
+Systems thus evaluated on private retrospective data wong2023scaling; yuan2023large; unlu2024retrieval cannot be transparently compared.kantor2024machine stress on standardized benchmarks as the dependence on manual evaluation hinders meta-analyses and comparison between different studies.
 Public annotations, such as the TREC CT tracks have an average of 700 trial annotations per patient for less than 100 patients, while the N2C2 has only a few thousands of criterion-level annotations.
-Jointly, the major public corpora, e.g., ClinicalTrials.gov and MIMIC datasets, present an opportunity to build on the proposal by Kim and Quintana (2022) to generate large-scale data using automated methods.
+Jointly, the major public corpora, e.g., ClinicalTrials.gov and MIMIC datasets, present an opportunity to build on the proposal by kim2022review to generate large-scale data using automated methods.
 Dimensions of Evaluation
-The benchmarks in Table 1 focus on model accuracy, corroborating the result from Bedi et al.
-(2025), which reports that more than 95% of studies use accuracy as the primary dimension of evaluation, while fairness, bias and uncertainty are less frequently measured.
-Omar et al.
-(2024) reviewed 27 clinical trials evaluating LLMs in healthcare also found the accuracy and reliability standards for LLM use to be undefined.
-Further results from Nemati et al.
-(2025), who benchmarked the annotation ability of LLMs across 9 performance metrics, show that while LLMs consistently score high on precision, recall and F1 (lowest being 0.8), their scores highly vary on semantic similarity, factual consistency, relevance, fluency, consistency and coherence (ranging from 0.1 to 0.9) highlighting the need for multiple dimensions of evaluation.
+The benchmarks in Table 1 focus on model accuracy, corroborating the result from bedi2025testing, which reports that more than 95% of studies use accuracy as the primary dimension of evaluation, while fairness, bias and uncertainty are less frequently measured.
+omar2024large reviewed 27 clinical trials evaluating LLMs in healthcare also found the accuracy and reliability standards for LLM use to be undefined.
+Further results from nemati2025benchmarking, who benchmarked the annotation ability of LLMs across 9 performance metrics, show that while LLMs consistently score high on precision, recall and F1 (lowest being 0.8), their scores highly vary on semantic similarity, factual consistency, relevance, fluency, consistency and coherence (ranging from 0.1 to 0.9) highlighting the need for multiple dimensions of evaluation.
 Formalization
@@
 Surprisingly, very few works explicitly formalize the matching task.
-This lack of formalization coincides with an absence of aggregation strategies for trial-level predictions Hamer et al.
-(2023); Wornow et al.
-(2025); Unlu et al.
-(2024) and others.
-Formalization guides Yuan et al.
-(2023) to design a loss function that accounts for the contrastive requirements of inclusion and exclusion criteria, Jullien et al.
-(2024) and Saeidi (2025) to define re-ranking scores.
+This lack of formalization coincides with an absence of aggregation strategies for trial-level predictions hamer2023improving; wornow2025zero; unlu2024retrieval and others.
+Formalization guides yuan2023large to design a loss function that accounts for the contrastive requirements of inclusion and exclusion criteria, jullien2024controlled and saeidi2025streamlining to define re-ranking scores.
 Societal Impact
-Recent research measured distinct bias in disease diagnosis across gender, age and disease in popular LLM models (GPT4, ChatGPT and Qwen) Zhao et al.
-(2024b) and found that GPT4 tends to stereotype demographic presentations when generating diagnoses Zack et al.
-(2024).
-Alber et al.
-(2025) show that LLMs are prone to making medical misjudgments by replacing just 0.001% of the training data with medical misinformation.
-All the LLM-based systems, except for the trial ranking models, are evaluated on a few disease-specific trials (see Table LABEL:tab:systems_overview_llm in Appendix C), the largest being 146 trials, covering 10 cancer types, evaluated by Hamer et al.
-(2023), making the generalizability of LLMs across diseases unclear.
+Recent research measured distinct bias in disease diagnosis across gender, age and disease in popular LLM models (GPT4, ChatGPT and Qwen) zhao2024can and found that GPT4 tends to stereotype demographic presentations when generating diagnoses zack2024assessing.
+alber2025medical show that LLMs are prone to making medical misjudgments by replacing just 0.001% of the training data with medical misinformation.
+All the LLM-based systems, except for the trial ranking models, are evaluated on a few disease-specific trials (see Table LABEL:tab:systems_overview_llm in Appendix C), the largest being 146 trials, covering 10 cancer types, evaluated by hamer2023improving, making the generalizability of LLMs across diseases unclear.
 Data Sensitivity
-While we found several methods that deployed GPT models on Azure AI to comply with privacy regulations Unlu et al.
-(2024); Wong et al.
-(2023); Wornow et al.
-(2025), processing patient data with LLMs still raises serious ethical challenges due to the lack of HIPAA compliance Edemekong et al.
-(2024).
+While we found several methods that deployed GPT models on Azure AI to comply with privacy regulations unlu2024retrieval; wong2023scaling; wornow2025zero, processing patient data with LLMs still raises serious ethical challenges due to the lack of HIPAA compliance edemekong2024health.
 As already discussed, creating large-scale realistic patient records while protecting their privacy is particularly challenging.
-There is interest in generating synthetic data, called digital twins, with limited access to real patient data as a viable privacy protected alternative Das et al.
-(2023); Wang et al.
-(2024).
+There is interest in generating synthetic data, called digital twins, with limited access to real patient data as a viable privacy protected alternative das2023twin; wang2024twin.
 Core Limitations of LLMs
-According to Harrer (2023) the core limitations affecting LLM adoption are unfiltered pre-training, which does not differentiate between facts, opinions, or misinformation; lack of self-assessment, where a model generates invalid but syntactically and semantically coherent sentences; non-determinism: where surface-form prompt variations lead to drastic changes in the output and repeatability is not guaranteed under consistent input conditions; and, knowledge recall: where updating outdated data or injecting new information requires expensive retraining since the mechanisms of memory in LLMs are not well understood.
+According to harrer2023attention the core limitations affecting LLM adoption are unfiltered pre-training, which does not differentiate between facts, opinions, or misinformation; lack of self-assessment, where a model generates invalid but syntactically and semantically coherent sentences; non-determinism: where surface-form prompt variations lead to drastic changes in the output and repeatability is not guaranteed under consistent input conditions; and, knowledge recall: where updating outdated data or injecting new information requires expensive retraining since the mechanisms of memory in LLMs are not well understood.
 In Section 7.1, we discussed how knowledge recall and lack of self-assessment surface through reasoning errors.
-These limitations pose a direct challenge to the transparency and accountability principles of AI for health laid down by the World Health Organization Guidance (2021).
+These limitations pose a direct challenge to the transparency and accountability principles of AI for health laid down by the World Health Organization guidance2021ethics.
 9 Future Directions
-As interest in LLMs orchestrating an end-to-end pipeline and incorporating human interactions is gaining more attention Gao et al. (2024); Qiu et al. (2024), we focus on four promising directions.
+As interest in LLMs orchestrating an end-to-end pipeline and incorporating human interactions is gaining more attention gao2024empowering; qiu2024llm, we focus on four promising directions.
 Trial Search
 Past trials that share a target population are important for designing new trials, recruiting patients, systematic reviews and meta-analyses.
-This problem has seen little activity since clustering using lexical features Hao et al.
-(2014).
-Newer search methods include constructing a clinical trial knowledge graph Chen et al.
-(2022), searching via patient EHRs Wu et al.
-(2018) and designing features for similarity matchingSun et al.
-(2022).
-Jullien et al.
-(2023) use textual entailment in LLMs to find trials that match short descriptions.
+This problem has seen little activity since clustering using lexical features hao2014clustering.
+Newer search methods include constructing a clinical trial knowledge graph chen2022knowledge, searching via patient EHRs wu2018semehr and designing features for similarity matchingsun2022interactivess.
+jullien2023nli4ct use textual entailment in LLMs to find trials that match short descriptions.
 However, these models falter on inferences that require numerical reasoning and could not surpass a BM25 baseline for ranking evidence.
@@
 LLM agents have the potential of bridging the semantic gap between eligibility criteria and patient data by suggesting data models underlying patient data for structuring eligibility criteria early on.
-This collaborative idea is not new Luo et al.
-(2013), yet, designing eligibility criteria is a big challenge and has been handled post-hoc by optionally considering patient data models Kang et al.
-(2017); Sun and Loparo (2019); Liu et al.
-(2021a); Dasgupta et al.
-(2020).
-Small patient pools, which ultimately affect the successful completion of a trial, are often the result of restrictive criteria Clinical Trials Arena (2022).
-LLMs can improve trial design by identifying restrictive criteria for trial investigators to relax them and create larger patient pools, especially for trials tackling diseases with a high mortality rate Liu et al.
-(2021b).
+This collaborative idea is not new luo2013human, yet, designing eligibility criteria is a big challenge and has been handled post-hoc by optionally considering patient data models kang2017eliie; sun2019knowledge; liu2021knowledge; dasgupta2020extracting.
+Small patient pools, which ultimately affect the successful completion of a trial, are often the result of restrictive criteria clinicaltrialarena2022.
+LLMs can improve trial design by identifying restrictive criteria for trial investigators to relax them and create larger patient pools, especially for trials tackling diseases with a high mortality rate liu2021evaluating.
 Collaborative Trial Planning
-Liu et al.
-(2025c) propose an iterative feature discovery model using LLM agents for interpretable trial outcome prediction.
-Markey et al.
-(2025) showed promising results on content relevance and suitability of trial protocols generated using LLMs, with room for improvement in logical reasoning and provenance.
-Similar to Shi et al.
-(2024), who propose collaboration of agents for knowledge-augmentation and reasoning, LLM agents can identify and distribute tasks and aggregate them to a final result.
-Another important operation is learning to defer to experts Mozannar and Sontag (2020), which can separate operable criteria, such as those that require tabular operations (e.g., via structured queries) or reasoning (temporal, numerical, negation), from criteria that require expert feedback.
+liu2025autoct propose an iterative feature discovery model using LLM agents for interpretable trial outcome prediction.
+markey2025rags showed promising results on content relevance and suitability of trial protocols generated using LLMs, with room for improvement in logical reasoning and provenance.
+Similar to shi2024enhancing, who propose collaboration of agents for knowledge-augmentation and reasoning, LLM agents can identify and distribute tasks and aggregate them to a final result.
+Another important operation is learning to defer to experts mozannar2020consistent, which can separate operable criteria, such as those that require tabular operations (e.g., via structured queries) or reasoning (temporal, numerical, negation), from criteria that require expert feedback.
 Explainable Matches
-Explaining black-box LLM predictions in human-understandable form is very challenging Zhao et al.
-(2024a).
-Explainability in clinical trial matches is limited to chain-of-thought generations, which is only one of the many facets of explainability Nauta et al.
-(2023); Bodria et al.
-(2023); Chen and Eickhoff (2024).
-Wong et al.
-(2023) could potentially explain eligibility prediction via logical component (mis)matches.
+Explaining black-box LLM predictions in human-understandable form is very challenging zhao2024explainability.
+Explainability in clinical trial matches is limited to chain-of-thought generations, which is only one of the many facets of explainability nauta2023anecdotal; bodria2023benchmarking; chen2024evaluating.
+wong2023scaling could potentially explain eligibility prediction via logical component (mis)matches.
 Furthermore, the utility of these explanations are unclear, as they have been only evaluated manually.
@@
 Appendix A Systematic Reviews versus Surveys
-A systematic review is a common practice in the medical research community, with standardized reporting guidelines Moher et al.
-(2009); Page et al.
-(2021).
+A systematic review is a common practice in the medical research community, with standardized reporting guidelines moher2009preferred; page2021prisma.
 This guideline has a checklist of required items in the title, abstract, introduction, methods, results, discussion and other information that every study needs to fulfil.
@@
 A survey is more flexible and provides a board coverage of the topic being discussed.
-For instance, the guidelines for writing surveys in our community as outlined by ACM Computing Surveys , TACL and ARR aim to draw perspectives on an evolving topic of interest.
+For instance, the guidelines for writing surveys in our community as outlined by ACMCSUR, TACL and ARR aim to draw perspectives on an evolving topic of interest.
 ACM Computing Surveys (long paper)
@@
 In the following we illustrate the the error distributions by percentages as reported by previous studies.
-Errors reported in Jin et al. (2024) supplementary:
+Errors reported in jin2024matching supplementary:
   • Incorrect reasoning: 30.7%
@@
   • other errors: 26.9% (self-conflicting)
-Errors reported in Hamer et al. (2023):
+Errors reported in hamer2023improving:
   • Incorrect reading: 6.5%
@@
   • Incorrect reasoning: 91.3%
-Errors reported in Nievas et al. (2024):
+Errors reported in nievas2024distilling:
   • Lack of knowledge  55%
@@
 We also notice a shift towards trial ranking due to publicly available data.
-Direction Method Patient Source Trial Source Limitations Trial-centric Penberthy et al. (2010) Discrete-data filter and sub-word matching Availability:
+Direction Method Patient Source Trial Source Limitations Trial-centric penberthy2010automated Discrete-data filter and sub-word matching Availability:
 Private
@@
 String similarity for keyword search.
-Yuan et al. (2019) (Criteria2Query)
+yuan2019criteria2query (Criteria2Query)
 NL to structured queries
@@
 Entity normalization into a small set of 2000 concepts.
-Tun et al. (2023)
+tun2023automatic
 Rule-based filters
@@
 Rules and classifiers are not robust to changes in criteria
-Ni et al. (2015)
+ni2015increasing
 Discrete-data filter
@@
 12.6% precision
-Patient-centric Ni et al. (2015)
+Patient-centric ni2015increasing
 Return top vector matches for a patient
 (same as previous row) 4% precision
-Zhang and Demner-Fushman (2017)
+zhang2017automated
 Bag-of-words feature vector
@@
 Task Method Patient Source Trial Source Limitations
-Criterion-level prediction Hamer et al. (2023)
+Criterion-level prediction hamer2023improving
 1-shot prompt to LLM for criteria level prediction with explanation
@@
 Majority of criterion-level errors due to incorrect reasoning (91%)
-Unlu et al. (2024)
+unlu2024retrieval
 RAG-based
@@
 RAG pipeline not published. Evaluated on a single trial.
-Ferber et al. (2024)
+ferber2024end
 Criterion-level boolean eligibility labels
@@
 Use of proprietary model limits evaluation on synthetic patient data.
-Beattie et al. (2024)
+beattie2024utilizing
 Vector database of patient notes. Prompt (criteria, top-k patient notes) pair to LLM
@@
 Expert guidance requires manual expertise for every criterion.
-Wornow et al. (2025)
+wornow2025zero
 Vector database of patient notes. Prompt (criteria, top-k patient notes) pair to LLM
@@
 67% incorrect decision despite correct reasoning.
-Saeidi (2025)
+saeidi2025streamlining
 Prompt LLM with a fine-tuned BERT-based concept embeddings from patient and criteria to predict binary eligibility labels.
 Availability: Public (2018 N2C2: Cohort Selection) LLM reasoning capabilities not explored.
-Trial-level prediction Wong et al. (2023)
+Trial-level prediction wong2023scaling
 LLM prompted with a template to generate structured forms of eligibility criteria
@@
 Trial-level performance is very low (F1 score range:[29.6-48])
-Yuan et al. (2023) Prompt LLMs to reformulate eligibility criteria Availability: Private
+yuan2023large Prompt LLMs to reformulate eligibility criteria Availability: Private
 Size: 825
@@
 High variance between different trials: F1 range: 0.39 - 0.96.
-Trial ranking Pradeep et al. (2022)
+Trial ranking pradeep2022neural
 Synthesis queries for initial trial retrieval.
@@
 Availability: Public (TREC CT 2021) Zero-shot relevance ranking is only slightly better than BM25. Fine-tuned models need re-training on new data or format.
-Zhuang et al. (2023)
+zhuang2023team
 Hybrid sparse-dense retriever for top-1000.
@@
 Validation on textual patient note does not transfer well to testing on structured patient note
-Cosine similarity between patient and trial embeddings obtained using GPT by Richmond and Deshpande (2023) and using Sentence Transformer by Lahiri et al.
-(2023).
-Availability: Public (TREC CT 2023) Very low MAP of 0.02 Richmond and Deshpande (2023) and close to zero for Lahiri et al.
-(2023)
-Kusa et al. (2023b)
-Sentence query formulation using GPT-3.5. Query enrichment using Kusa et al. (2023a). Trial-level prediction of re-ranked pairs using GPT-3.5.
+Cosine similarity between patient and trial embeddings obtained using GPT by richmond2023leveraging and using Sentence Transformer by lahiri2023tmu.
+Availability: Public (TREC CT 2023) Very low MAP of 0.02 richmond2023leveraging and close to zero for lahiri2023tmu
+kusa2023dossier
+Sentence query formulation using GPT-3.5. Query enrichment using kusa2023effective. Trial-level prediction of re-ranked pairs using GPT-3.5.
 Availability: Public (TREC CT 2023) Zero-shot LLM prompts provide marginal improvement over neural-rerankers
-Peikos (2023)
+piekos2023unimib
 Utilize GPT3.5 to obtain trial-level labels for final re-ranking on top of lexical and neural re-rankers
 Availability: Public (TREC CT 2023) Query processing is template-based and excludes negative information.
-Ferber et al. (2024)
+ferber2024end
 No-SQL query formulation from patient EHR using GPT-4o for initial retrieval, followed by vector embedding re-ranking.
@@
 Use of proprietary model limits evaluation on synthetic patient data.
-Rybinski and Karimi (2023); Rybinski et al. (2024)
+rybinski2023matching; rybinski2024learning
 Multi-stage retriever using LLMs
 Availability: Public (TREC CT 2021, 2022, 2023) High latency of GPT4 vs a small performance boost
-Jullien et al. (2024)
+jullien2024controlled
 LLM-guided basic attribute extraction for trial retrieval and filtering. Criterion-level LLM predictions fed to set-reasoning-based re-rank scoring functions.
 Availability: Public (TREC CT 2022) LLMs underperform in exclusion criteria labelling.
-Jin et al. (2024)
-Hybrid trial filtering with BM25 and MedCPT Jin et al. (2023)
+jin2024matching
+Hybrid trial filtering with BM25 and MedCPT jin2023medcpt
 Criterion-level LLM prediction aggregated to trial scores
-Availability: Public (Koopman and Zuccon (2016), TREC CT 2021, 2022)
+Availability: Public (koopman2016test, TREC CT 2021, 2022)
 Additional annotation: 1,015 criterion-patient labels
@@
 LLM aggregation requires LLM to perform mathematical reasoning and computations
-Nievas et al. (2024)
-Extends Jin et al. (2024) to open-source LLMs
-(same as Jin et al. (2024))
+nievas2024distilling
+Extends jin2024matching to open-source LLMs
+(same as jin2024matching)
 Additional annotation: Patient sentence supporting eligibility label Additional annotation: 500 criteria labels on: eligibility and difficulty.
 Significantly high fine-tuning costs
-Datta et al. (2025)
+datta2025patient2trial
 Lexical retrievers use LLM generated query expansion. LLM predicts trial-level label and a criterion-level rationale with a matching score. Final ranking based on matching score.
@@
 Model predicted trial-level label is not evaluated
-Saeidi et al.
-(2023); Saeidi (2025) Embed patients and trials to concept vector space.
+saeidi2023malnis; saeidi2025streamlining Embed patients and trials to concept vector space.
 Use variations of Equation 6.1 to compute relevance scores.